### v1

# MacRAG: Compress, Slice, and Scale-up for Multi-Scale Adaptive Context RAG

Abstract
Long-context (LC) Large Language Models (LLMs) combined with Retrieval-Augmented Generation (RAG) hold strong potential for complex multi-hop and large-document tasks. However, existing RAG systems often suffer from imprecise retrieval, incomplete context coverage under constrained context windows, and fragmented information caused by suboptimal context construction. We introduce Multi-scale Adaptive Context RAG (MacRAG), a hierarchical retrieval framework that compresses and partitions documents into coarse-to-fine granularities, then adaptively merges relevant contexts through chunk- and document-level expansions in real time. By starting from the finest-level retrieval and progressively incorporating higher-level and broader context, MacRAG constructs effective query-specific long contexts, optimizing both precision and coverage. Evaluations on the challenging LongBench expansions of HotpotQA, 2WikiMultihopQA, and Musique confirm that MacRAG consistently surpasses baseline RAG pipelines on single- and multi-step generation with Llama-3.1-8B, Gemini-1.5-pro, and GPT-4o. Our results establish MacRAG as an efficient, scalable solution for real-world long-context, multi-hop reasoning. Our code is available at https://github.com/Leezekun/MacRAG.
MacRAG: Compress, Slice, and Scale-up
for Multi-Scale Adaptive Context RAG
Woosang Lim1thanks: Equal contribution.thanks: Corresponding authors: woosang.quasar@gmail.com, william@cs.ucsb.eduthanks: Work done while at POSCO HOLDINGS.   Zekun Li211footnotemark: 1   Gyuwan Kim211footnotemark: 1   Sungyoung Ji [^1] footnotemark: 1 HyeonJung Kim1   Kyuri Choi1   Jin Hyuk Lim1   Kyungpyo Park3   William Yang Wang [^2] footnotemark: 2 POSCO HOLDINGS1   University of California, Santa Barbara2   Google Cloud3
1 Introduction
Figure 1: An overview of the MacRAG framework, consisting of two main phases: (1) top-down hierarchical indexing (upper), and (2) bottom-up multi-scale adaptive retrieval on the constructed hierarchy of document-chunk-summary-slice (lower).
Large Language Models (LLMs) have substantially advanced question answering and reasoning, yet they frequently exhibit factual gaps or hallucinations when relying solely on internal parameters. Retrieval-Augmented Generation (RAG) (Guu et al., 2020; Lewis et al., 2020) mitigates these pitfalls by equipping LLMs with external evidence from large document collections. Meanwhile, Long Context (LC) LLMs, such as GPT-4o (OpenAI, 2024), Gemini 1.5 (Team et al., 2024), and Llama 3 (Dubey et al., 2024a), push context windows to hundreds of thousands or even millions of tokens, enabling an entire, particularly long context or document to be processed in principle. However, even these extended windows do not fully overcome the “Lost in the Middle” phenomenon (Liu et al., 2024; Xu et al., 2024b), wherein crucial mid-text information is overlooked, and their performance may degrade once length thresholds are surpassed (Leng et al., 2024; Yu et al., 2024).
Within RAG, these issues manifest as three core trade-offs in document processing and retrieval for answer generation:
  1. (1)
    Context Length vs. Focus. Long contexts can improve recall but risk eclipsing critical details (particularly mid-text), whereas shorter contexts narrow in on precision but risk omitting essential evidence.
  2. (2)
    Chunking & Indexing. Finer-grained chunks support precise retrieval but can fragment continuity, while larger chunks preserve coherence but become redundant or unwieldy.
  3. (3)
    Coverage vs. Computation. Contexts across documents for multi-relational queries strain token budgets, latency limits or expensive pre-processing/updates, forcing a balance between broad coverage and efficiency on inference and system operations
These trade-offs become especially acute in enterprise-grade or domain-specific deployments that demand “almost infinite” context for personalization or massive corpora. Moreover, recent iterative RAG approaches Yue et al. (2024); Wang et al. (2025) and agentic RAG pipelines Asai et al. (2024); Jeong et al. (2024); Chen et al. (2025) repeatedly invoke retrieval and generation, compounding inefficiencies unless the underlying retrieval module is robust.
To systematically address these trade-offs, we propose Multi-scale Adaptive Context RAG (MacRAG), a hierarchical framework that unifies top-down offline indexing with bottom-up query-time retrieval. In an offline phase, MacRAG partitions each document into partially overlapping chunks, compresses them (e.g., via abstractive summarization), and further slices these compressed segments for fine-grained indexing via standard overlap chunking. This creates a hierarchical index from document level to fine-grained slices. At query time, MacRAG retrieves the most relevant slices at the finest level to ensure high precision, then progressively scales up to find more borderline contexts. It merges neighboring chunks and performs document-level expansions to construct effective long contexts in real time while controlling context size, addressing precision, coverage, and efficiency trade-offs simultaneously. This hierarchical multi-scale approach is compatible with both single- and multi-step generation methods (Xu et al., 2024a; Zhao et al., 2024) and integrates readily into iterative or agent-based RAG systems. Unlike prior works that improve either retrieval or ranking alone, MacRAG integrates structure-preserving indexing, bottom-up adaptive retrieval, and bounded context assembly for robust multi-hop reasoning. Through extensive evaluations on diverse long-document collections and multi-hop reasoning tasks, MacRAG demonstrates substantial improvements across various generation methods when used as a core RAG module with Llama-3.1-8B, Gemini-1.5-pro, and GPT-4o.
2 Related Work
2.1 Multi-step Generation with Long Context in RAG
Despite advances in LC LLMs, their integration with RAG systems continues to face the first trade-off discussed in Section 1. Recent research has explored context compression solutions to address the “Lost in the Middle" phenomenon and diminishing returns from excessive context.
Recent approaches have explored two main techniques to effectively handle long contexts via multi-step generation with compression and filtering. Abstractive summarization methods, such as RECOMP-Abst Xu et al. (2024a), generate condensed text by rewriting retrieved passages. Extractive summarization approaches, including RECOMP-Extr Xu et al. (2024a), LLMLingua Jiang et al. (2023a, b), and EXIT Hwang et al. (2024), preserve original text through selective extraction with verification. LongRAG Zhao et al. (2024) offers a different perspective by exploring multiple strategies for utilizing retrieved chunks, comparing direct chunk usage against whole document incorporation and introducing a chain-of-thought-based filter and extractor for selecting relevant information.
However, these post-retrieval multi-step generation approaches for long context face fundamental limitations. First, they rely exclusively on post-retrieval context compression without improving initial retrieval quality, creating an inherent performance ceiling. Second, generation based on compressed rather than original context risks information loss. Third, they incur significant computational overhead through multi-step processing on potentially extensive long context input, i.e., from corpora consisting of excessively long documents. In contrast, MacRAG focuses on the retrieval phase, efficiently constructing precise and comprehensive context inputs with bounded length for generation while minimizing related information loss.
2.2 Hierarchical Retrieval Approaches for RAG
Hierarchical retrieval approaches offer an alternative to post-retrieval compression by capturing longer or broader contexts during the retrieval phase. These methods organize information in hierarchical structures, aiming to improve both precision and recall.
Sarthi et al. (2024) introduce RAPTOR, a tree-based architecture that recursively summarizes clustered text chunks for semantic matching. However, its exclusive focus on semantic similarity presents two key limitations: (1) it often misses important connected relationships that lack direct semantic similarity, hindering multi-hop reasoning capabilities, and (2) its rigid hierarchical structure can fragment interconnected knowledge. SIRERAG Zhang et al. (2025) extends this approach by incorporating relational connectivity alongside semantic similarity in a dual-tree framework. Although this enables more comprehensive knowledge synthesis, it introduces significant computational overhead and depends heavily on accurate proposition extraction and entity-based clustering. The system may miss subtle connections that fall outside its predefined relational model.
Both approaches face substantial computational and practical limitations. Their clustering-based strategies incur high costs during database updates, which is particularly challenging for dynamic knowledge bases. Furthermore, these methods struggle with information overlap across hierarchical levels and lack robust mechanisms to balance semantic similarity with relational relevance. In contrast, MacRAG is more flexible and efficient by preserving the original content structure during indexing and retrieval.
Recent work, such as GraphRAG (Edge et al., 2024) and HippoRAG (Gutiérrez et al., 2024), adopts hierarchical retrieval through symbolic graphs or sentence-level indexing, which improves coverage but introduces additional overhead. In contrast, MacRAG achieves similar benefits via dense, bottom-up expansion over pre-indexed multi-scale units without symbolic structure. Another line of work bypasses chunking altogether by retrieving entire passages or documents (Qian et al., 2024). While this simplifies retrieval, it reduces granularity control. MacRAG retains a flexible structure through hierarchical slicing and adaptive composition.
3 Multi-Scale Adaptive Context RAG
We propose Multi-Scale Adaptive Context RAG (MacRAG), a hierarchical multi-scale adaptive retrieval system composed of three main components in sequence. First, the top-down hierarchical indexing processes documents via compression, chunking, and slicing, building a hierarchical index from coarse-grained documents to fine-grained slices. Second, the bottom-up multi-scale adaptive retrieval begins at the finest granularity and progressively traces back to broader contexts, merging relevant information to construct query-adaptive long contexts in real-time. Third, the response generation leverages these merged contexts to produce the final output.
Our primary contribution lies in the retrieval pipeline (phases 1 and 2), which offers a modular foundation for integrating arbitrary response generation techniques in phase 3. Figure 1 concisely illustrates MacRAG, while Algorithm 1 describes its implementation, with mathematical details elaborated in the following sections. The retrieval process is adaptive in the sense that the amount and granularity of retrieved content vary depending on the query’s ambiguity and the initial slice-level relevance scores.
3.1 Document Hierarchical Indexing
Given a document corpus $\mathcal{D}=\{d_{1},\dots,d_{N}\}$, MacRAG performs offline processing to construct a hierarchy through chunking (§3.1.1), compression (§3.1.2), and slicing (§3.1.3), while building a slice-level vector DB at the finest granularity to enable retrieval. Instead of using raw sentences as the base unit, we use partial summaries to reduce redundancy and enhance semantic focus, which helps improve retrieval precision and reduces noise in the multi-scale expansion.
Algorithm 1 MacRAG: Multi-Scale Adaptive Context RAG
Query $q$, document corpus $\mathcal{D}$, $k_{1}$ (for slice retrieval), $k_{2}$ (for final merged chunks), the number of hops $h$, up-scaling factor $\alpha$
Top-$k_{2}$ merged chunks $\mathcal{C}_{\mathrm{final}}$, generated answer
Phase 1: Top-down Hierarchical Indexing
For each document $d\in\mathcal{D}$:
(1.1) Split $d$ into a set of overlapping chunks $\mathcal{C}_{d}$.
(1.2) Compress each chunk $c_{i}$ into a shorter summary $s_{i}$ via summarization.
(1.3) Split each compressed summary $s_{i}$ into overlapping slices $\mathcal{L}_{i}$.
(1.4) Encode each slice and store its embedding & metadata (doc ID, chunk ID, offset, etc) in the database.
Phase 2: Bottom-up Multi-Scale Adaptive Retrieval
(2.1) Retrieve the top-$k_{1}$ slices based on similarity to $q$.
(2.2) Map retrieved slices to their parent chunks to obtain $\mathcal{C}_{k_{1}}^{\prime}$.
(2.3) Rerank chunks in $\mathcal{C}_{k_{1}}^{\prime}$ to refine their scores $r^{\prime}_{q,c_{i}}$.
(2.4) Scale-up by selecting the top $(k_{2}\times\alpha)$ chunks and ranking their source documents by aggregated scores to choose $k_{2}$ distinct documents.
(2.5) Merge the top-ranked chunks’ $h$-hop neighbors to form the final top-$k_{2}$ chunks, $\mathcal{C}_{\mathrm{final}}$.
Phase 3: Response Generation
Provide $\mathcal{C}_{\mathrm{final}}$ as context to an LLM (or any downstream model) to generate the final answer.
3.1.1 Chunking
Each document $d\in\mathcal{D}$ is split into partially overlapping chunks,
$\mathcal{C}_{d}=\{c_{1},c_{2},\dots,c_{m}\},$ (1)
where each chunk $c_{i}$ spans approximately 200 to 500 tokens, with a fixed overlap of 10–20 tokens between consecutive chunks. This controlled overlap preserves semantic continuity across chunk boundaries. Although the chunking is token-based by default, it can alternatively be configured at the character level.
3.1.2 Compression
Each chunk $c_{i}$ is transformed into a more compressed summary $s_{i}$ and thus we obtain a set of compressed summaries $\mathcal{S}_{d}=\{\,s_{1},\dots,s_{m}\}$, where $s_{i}=\mathrm{Compress}(c_{i})$ and $\mathrm{Compress}(\cdot)$ represents a summarization model. It can be either extractive (selecting salient sentences) or abstractive (using a generative model to generate a summary). By default, we use an LLM to generate the abstractive summarization. This step reduces redundancy while retaining core factual and named-entity information.
3.1.3 Slicing
To facilitate finer-grained retrieval, each summary $s_{i}$ is further split into overlapping slices: $\mathcal{L}_{i}=\bigl{\{}\,\ell_{i,1},\ell_{i,2},\dots,\ell_{i,j},\dots\bigr{\}}$, where each slice $\ell_{i,j}$ typically spans $50\text{-}200$ tokens with partial overlap as in the standard overlap chunking. Each slice is then embedded via $\mathbf{e}_{i,j}=\mathrm{Embed}(\ell_{i,j})$, where $\mathrm{Embed}(\cdot)$ represents a text encoder (e.g., multilingual-e5-large). The slice-level embeddings with metadata (document ID, chunk ID, offsets) are indexed in the vector DB. MacRAG can optionally index chunk and summary embeddings for multi-view retrieval.
3.2 Multi-Scale Adaptive Retrieval
Given an input query $q$, MacRAG implements a multi-scale adaptive retrieval pipeline over the constructed hierarchy through five sequential steps, ultimately selecting the top-$k_{2}$ merged chunks with contextually relevant neighbor expansions for enhanced retrieval precision and recall.
3.2.1 Initial Slice-level Retrieval
With slices serving as the finest granular units in our hierarchical system and vector DB, we first compute the query-slice relevance scores for each slice $\ell_{i,j}$ in the vector DB:
$r_{q,\ell_{i,j}}=\mathrm{Rel}\!\bigl{(}q,\,\ell_{i,j}\bigr{)},$ (2)
where $\mathrm{Rel}(\cdot,\cdot)$ can be implemented with relevance model, e.g., cosine similarity and $L_{2}$ distance in dense retrieval, sparse retrieval, hybrid retrieval, and so on. We then retrieve the top-$k_{1}$ most relevant slices. This fine-grained retrieval ensures high precision by identifying slices with strong semantic alignment to query $q$.
3.2.2 Parent Chunk Mapping
Each retrieved slice corresponds to a parent chunk $c_{i}$. To prevent redundancy when multiple retrieved slices originate from the same chunk, we perform a unique mapping:
$\mathcal{C}_{k_{1}}^{{}^{\prime}}=\mathrm{unique}\bigl{(}{c_{i}\mid\ell_{i,\cdot}\in\mathrm{top}\text{-}k_{1}(r_{q,\ell})}\bigr{)},$ (3)
where we identify and consolidate the parent chunks of all retrieved slices and obtain a set of selected chunks used for the next steps.
3.2.3 Chunk-level Re-Ranking
For each identified unique relevant chunk from the mapping as $c_{i}\in\mathcal{C}_{k_{1}}^{{}^{\prime}}$ to the query, we compute a chunk-level ranking score for the query:
$r_{q,c_{i}}=\mathrm{ReRank}\bigl{(}q,\,c_{i}\bigr{)},$ (4)
utilizing a cross-encoder (e.g., marco-miniLM) or advanced ranking model that evaluates the complete chunk content, mitigating potential “information fragmentation" inherent in slice-level retrieval. The chunks in $\mathcal{C}_{k_{1}}^{{}^{\prime}}$ are then reordered according to their ranking scores to the query $q$.
3.2.4 Scaling-Up & Document-level Ranking
In this step, we aim to select content segments, which can be used to construct $k_{2}$ precise long contexts with upper bounds in the later step to ensure both relevance and sufficient context coverage for final response generation with LLM. To achieve this, we incorporate top-ranked chunks and their document neighbors, thereby mitigating information fragmentation and enabling multi-hop reasoning. The process consists of the following steps.
Scaled Top Chunk Selection:  Instead of picking exact top-$k_{2}$ chunks, we consider more borderline chunk candidates from the re-ranking stage, since the computed $r_{q,c_{i}}$ may not be perfect. Thus, MacRAG adopts scaling factor $\alpha$ and selects the top-($k_{2}\times\alpha$) chunks from $\mathcal{C}_{k_{1}}^{{}^{\prime}}$ for up-scaling with $r_{q,c_{i}}$ in Equation 4:
[…]
where $\alpha\geq 1$ and $(k_{2}\times\alpha)\leq|\mathcal{C}^{{}^{\prime}}_{k_{1}}|$. $\alpha=1$ does not scale-up the number of borderline candidates, but moderate higher $\alpha$ (e.g., 3 or 4) can provide higher probabilities to include borderline but crucial chunks to improve coverage while maintaining certain-level of qualities for candidates. To balance among precision, recall, and context lengths, $\alpha\in\{2,3,4\}$ would be promising where $(k_{2}\times\alpha)\leq|\mathcal{C}^{{}^{\prime}}_{k_{1}}|$, and we will also demonstrate that selection $\alpha$ is not sensitive, but robust with handling trade-off problems well in Section 4. This approach can also potentially facilitate bridging multi-hop queries.
Document Selection:  We further track one level up from the $(k_{2}\times\alpha)$ chunks $c_{i}$ in $\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$ by mapping $c_{i}$ into their parents to identify unique candidates for selecting top-$k_{2}$ documents. We then compute the rank score $r_{q,d}$ of mapped document with re-ranking mechanism leveraging chunks’ rank score $r_{q,c_{i}}$ from all the chunks $c_{i}$ within the document $d$ and $c_{i}$ in $\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$:
$r_{q,d}\;=\;\mathrm{ReRank_{Doc}}\Bigl{(}\{c_{i}\in d~{}and~{}c_{i}\in\mathcal{C}^{{}^{\prime}}_{q,k_{2}}\}\Bigr{)}.$ (6)
Potential choices for document ranking functions could be mean, max, sum, or an LLM-based ranker. If chunk’s rank score $r_{q,c_{i}}$ is computed by using a cross encoder, then it can be either positive and negative, then $\max({r_{q,c_{i}}})$ would be the base function for document ranking. Based on the computed $r_{q,d}$, we construct the $\mathcal{D}_{q,k_{2}}$ consisting of top-$k_{2}$ documents. We ensure top-$k_{2}$ distinct docs so coverage is not restricted to a single or few docs. If we consider properly higher $\alpha$ (e.g., 3 or 4 depending on the overlapped ratio of original chunks), then there are fewer risks for unfilled top-ranked documents. This is critical when relevant evidence is scattered across multiple documents.
3.2.5 Neighbor Chunk Propagation & Merging
Within each of the top-$k_{2}$ distinct documents, we focus on chunks $c_{i}$ both in $\mathcal{D}_{q,k_{2}}$ and $\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$, as these represent the most relevant content segments. For such a chunk $c_{i}$, we define $\mathcal{N}_{h}(c_{i})$ which consists of $c_{i}$’s neighbors within $h$-hops and $c_{i}$ itself, and process extension of $c_{i}$ to get $\mathcal{N}(c_{i})$ by using its associated indices,
$\mathcal{N}_{h}(c_{i})=\;\bigl{\{}\,c_{i-h},\dots,c_{i},\dots,c_{i+h}\bigr{\}}.$ (7)
We then merge $\mathcal{N}_{h}(c_{i})$ within the same ranked document to form the final merged chunk for adaptive long context to the query $q$:
$c^{merge}_{q,d}~{}=~{}$ $\mathrm{Merge}\bigl{(}\mathcal{N}_{h}(c_{i})\bigr{)},$ (8)
$\text{where}~{}d\in$ $\mathcal{D}_{q,k_{2}},~{}c_{i}\in d~{}and~{}c_{i}\in\mathcal{C}^{{}^{\prime}}_{q,k_{2}}.$
As described in Equation 8, from the top-$k_{2}$ documents in $\mathcal{D}_{q,k_{2}}$, we can select top-$k_{2}$ merged chunks $c^{merge}_{q,d}$ by considering its order along with top document order with $r_{q,d}$. Then, finally we have top-$k_{2}$ merged chunks as
$\mathcal{C}_{\mathrm{final}}~{}=~{}\mathrm{TopK}_{k_{2}}\!\Bigl{(}c^{merge}_{q,d}\Bigr{)}.$ (9)
The final merged chunks $\mathcal{C}_{\text{final}}$ provide diverse views across the top-$k_{2}$ documents while optimizing context length (Zhao et al., 2024), combining the most relevant content with minimal redundancy.
Mode Description
R&B Single-step generation using retrieved top-$k_{2}$ chunks, which is the vanilla basic version.
R&L Single-step generation using complete documents from top-$k_{2}$ chunks. Prone to “Lost in the Middle” effect and inefficient for long documents.
Full_Ext Multi-step generation: (1) LLM extracts query-relevant content from full documents of top-$k_{2}$ chunks, (2) Generates answer using both chunks and extracted content.
Fil Filtering re-ranked top-$k_{2}$ chunks, e.g., via LLM. Improves precision but may lose borderline info.
Full_E&F Multi-step generation combining Full_Ext’s extracted content and Fil’s filtered chunks. Balances coverage and precision, but expensive for long doc.
R&B_Ext Multi-step generation with content extraction limited to top-$k_{2}$ chunks, reducing computational overhead.
R&B_E&F Multi-step generation combining R&B_Ext’s extracted content and Fil’s filtered chunks. Optimizes precision-recall trade-off w/o processing full doc.
Table 1: Seven different generation methods with single-step and multi-step generation schemes using retrieved top-$k_{2}$ chunks.
3.3 Single/Multi-step Post Retrieval Generation
Our retrieval framework is modular, enabling integration with various post-retrieval generation approaches. These include both vanilla single-step generation directly using the retrieved top-$k_{2}$ chunks in a single pass and iterative multi-step generation as in LongRAG (Zhao et al., 2024), which introduces CoT-guided filter (which checks chunk-level relevance) and LLM-augmented extractor (which locates global context in original passages). Following the comprehensive exploration of single and multi-step generation variants in LongRAG, we evaluate MacRAG’s performance when combined with different post-retrieval generation methods. Specifically, we investigate seven generation methods, as detailed in Table 1. Five of these methods—R&B, R&L, Full_Ext, Fil, and Full_E&F—are adopted directly from LongRAG. Among these, R&L, Full_Ext, and Full_E&F process the entire set of top-$k_{2}$ documents, which can introduce computational overhead when handling very long texts. To address this limitation while maintaining an optimal balance between precision and recall, we introduce two new variants: R&B_Ext and R&B_E&F, to reduce this overhead while maintaining precision-recall balance.
3.4 Implementation Complexity and System Efficiency
MacRAG enhances simple overlapping-chunk indexing by first chunking each document with partial overlaps, then compressing each chunk through summarization, and finally slicing it. These processes are performed once during offline preparation, eliminating query-time overhead. Moreover, this document-level independence simplifies updates and streamlines database operations, avoiding the complexities associated with hierarchical multi-clustering Sarthi et al. (2024); Zhang et al. (2025).
At inference time, MacRAG constructs a real-time effective long context by mapping slices back to their parent chunks and documents using lightweight index-based lookups rather than costly tree traversals. This design leverages metadata dictionaries (document ID, chunk ID, token offsets) to enable modular neighbor merging with minimal overhead. A bounded context size is enforced through parameters $(k_{2}\times\alpha)$ and chunk dimensions, ensuring scalability to enterprise-level corpora without degrading performance. This controlled scalability, along with real-time context assembly, makes MacRAG highly valuable for massive long-document collections, “almost infinite context” in advanced personalization, and iterative and agentic RAGs.
Model HotpotQA 2WikimultihopQA Musique Average
Long Context (LC) LLM without RAG
Gemini-1.5-pro Team et al. (2024) 36.79 38.31 20.09 31.73
GPT-4o OpenAI (2024) 46.76 40.62 30.76 39.38
Agentic RAG Method
CRAG (GPT-3.5-Turbo) Yan et al. (2024) 52.04 41.13 25.34 39.50
Self-RAG (GPT-3.5-Turbo) Asai et al. (2024) 50.51 46.75 24.62 40.63
RAG with Reranking (R&B, Base Version)
Llama3-8B-8k Dubey et al. (2024b) 48.25 43.47 19.66 37.13
GPT-3.5-Turbo 52.31 43.44 25.22 40.32
Llama-3.1-8B-instruct (abbr. Llama-3.1-8B) 52.50 46.33 26.70 41.84
RAPTOR Sarthi et al. (2024) (Llama-3.1-8B) 52.30 (-0.20$\downarrow$) 43.61 (-2.72$\downarrow$) 23.79 (-2.91$\downarrow$) 39.90 (-1.94$\downarrow$, -4.64$\downarrow$)
MacRAG (Ours) (Llama-3.1-8B) 57.39 (+4.89$\uparrow$) 44.87 (-1.46$\downarrow$) 30.38 (+3.68$\uparrow$) 44.21 (+2.37$\uparrow$, +5.66$\uparrow$)
LongRAG Zhao et al. (2024) (Multi-step Generation, Full_E&F in Table 1)
LongRAG (GPT-3.5-Turbo) 56.17 51.37 32.83 46.79
LongRAG (Llama-3.1-8B) 58.49 55.90 32.73 49.04
MacRAG (Ours) (Llama-3.1-8B) 61.10 (+2.61$\uparrow$) 57.74 (+1.84$\uparrow$) 34.43 (+1.70$\uparrow$) 51.09 (+2.05$\uparrow$, +4.18$\uparrow$)
LongRAG (Gemini-1.5-pro) 61.95 58.17 34.04 51.39
MacRAG (Ours) (Gemini-1.5-pro) 64.39 (+2.44$\uparrow$) 64.75 (+6.58$\uparrow$) 43.34 (+9.30$\uparrow$) 57.49 (+6.10$\uparrow$, +11.87$\uparrow$)
LongRAG (GPT-4o) 66.20 65.89 43.83 58.64
MacRAG (Ours) (GPT-4o) 68.52 (+2.32$\uparrow$) 73.19 (+7.30$\uparrow$) 50.09 (+6.26$\uparrow$) 63.93 (+5.29$\uparrow$, +9.02$\uparrow$)
Table 2: Experimental results (F1-score) comparing RAPTOR, LongRAG and MacRAG on HotpotQA, 2WikimultihopQA, and MuSiQue datasets from LongBench Bai et al. (2023). Gains are displayed in parentheses with absolute number and its relative percentage.
Dataset Avg. Performance on LongRAG MacRAG LongRAG MacRAG
Seven Gen. Settings (Gemini-1.5-pro) (Gemini-1.5-pro) (GPT-4o) (GPT-4o)
HotpotQA Exact Match 46.48 48.36 (+1.88$\uparrow$, +4.04$\uparrow$) 51.85 53.60 (+1.75$\uparrow$, +3.37$\uparrow$)
F1-score 61.45 63.97 (+2.52$\uparrow$, +4.10$\uparrow$) 66.17 67.51 (+1.34$\uparrow$, +2.03$\uparrow$)
Precision 65.93 67.98 (+2.05$\uparrow$, +3.11$\uparrow$) 68.66 70.10 (+1.44$\uparrow$, +2.10$\uparrow$)
Recall 61.59 64.47 (+2.88$\uparrow$, +4.68$\uparrow$) 68.10 69.36 (+1.26$\uparrow$, +1.85$\uparrow$)
2WikimultihopQA Exact Match 49.93 53.69 (+3.76$\uparrow$, +7.53$\uparrow$) 52.90 57.60 (+4.69$\uparrow$, +8.87$\uparrow$)
F1-score 57.73 61.97 (+4.24$\uparrow$, +7.34$\uparrow$) 62.69 67.45 (+4.76$\uparrow$, +7.59$\uparrow$)
Precision 57.97 61.69 (+3.72$\uparrow$, +6.41$\uparrow$) 62.22 66.48 (+4.26$\uparrow$, +6.85$\uparrow$)
Recall 60.17 65.10 (+4.93$\uparrow$, +8.20$\uparrow$) 66.47 72.00 (+5.53$\uparrow$, +8.32$\uparrow$)
Musique Exact Match 26.05 33.75 (+7.70$\uparrow$, +29.57$\uparrow$) 31.30 36.48 (+5.18$\uparrow$, +16.55$\uparrow$)
F1-score 33.92 42.84 (+8.93$\uparrow$, +26.34$\uparrow$) 41.43 47.91 (+6.48$\uparrow$, +15.64$\uparrow$)
Precision 34.69 43.29 (+8.60$\uparrow$, +24.78$\uparrow$) 40.80 47.07 (+6.27$\uparrow$, +15.37$\uparrow$)
Recall 35.69 45.16 (+9.46$\uparrow$, +26.50$\uparrow$) 45.06 51.67 (+6.61$\uparrow$, +14.67$\uparrow$)
Table 3: Consolidated results for four performance metrics across three datasets, comparing LongRAG and MacRAG using the Gemini-1.5-pro and GPT-4o models. The results are averaged over seven generation settings in Table 1.
Figure 2: Performance comparison of GPT-4o using LongRAG and MacRAG across the seven settings in Table 1, showing F1-scores for three datasets (HotpotQA, 2WikiMultihopQA, and MuSiQue). Complete results for all metrics and LLMs are provided in the Appendix.
4 Experiments
4.1 Experimental Setup
Benchmark Datasets and Challenges:
We evaluate MacRAG on three multi-hop question answering benchmarks from the expanded LongBench framework (Bai et al., 2023): HotpotQA, 2WikimultihopQA, and Musique. These datasets have become the standard evaluation suite for hierarchical retrieval research, including recent systems such as SIRERAG Zhang et al. (2025), RAPTOR Sarthi et al. (2024), and LongRAG Zhao et al. (2024). Their LongBench variants specifically intensify the “Lost in the Middle" phenomenon by introducing near-duplicate and tangential passages, creating challenging retrieval environments where critical connecting information becomes obscured by surrounding context. Each dataset introduces distinct multi-hop reasoning challenges that systematically test different aspects of retrieval capability: HotpotQA features bridging and comparison queries requiring evidence synthesis across multiple paragraphs; 2WikiMultihopQA expands partial excerpts to complete articles, significantly increasing the risk that essential connecting information becomes buried within expansive context; and Musique embeds nested sub-questions within tangential sections, challenging systems to identify and connect scattered bridging facts across document boundaries. These characteristics directly test a retrieval system’s ability to handle fragmented evidence, reconcile overlapping information, and identify critical bridging facts despite contextual noise (Yue et al., 2024; Leng et al., 2024), precisely the challenges that MacRAG’s multi-scale architecture addresses through hierarchical indexing and adaptive retrieval.
Models and Evaluation Metrics:
Based on SIRERAG’s comparative analysis (Table 4 in Zhang et al. (2025)), we focus our evaluation on RAPTOR Sarthi et al. (2024) and LongRAG Zhao et al. (2024) as representative state-of-the-art hierarchical RAG systems. SIRERAG’s analysis establishes that GraphRAG underperforms on these specific multi-hop tasks, while RAPTOR, HippoRAG, and SIRERAG achieve comparable performance in terms of F1-score on multiple multi-hop question answering datasets Zhang et al. (2025). Among these top-performing systems, we selected RAPTOR and LongRAG due to their better balance between performance and computational efficiency, offering lower computational costs and easier maintenance compared to SIRERAG’s higher overhead.
We conducted experiments using three prominent long-context LLMs: two widely adopted proprietary models (GPT-4o and Gemini-1.5-pro) and one popular open-source model (Llama-3.1-8B-instruct). This diverse selection enables assessment of MacRAG’s generalizability across different model architectures and parameter scales.
We measure performance through four complementary metrics: Exact Match, F1-score, Precision, and Recall, providing a multidimensional view of answer quality. We additionally track average context length to quantify MacRAG’s efficiency in reducing input redundancy while maintaining or improving information coverage.
To ensure fair comparison across all methods (MacRAG, LongRAG, and RAPTOR), we basically applied the same re-ranking model (marco-miniLM) for all RAG methods.
Comparative Framework and Hyperparameters:
To ensure rigorous comparative analysis across all methods (MacRAG, LongRAG, and RAPTOR), we used the same re-ranking model (marco-miniLM) for all RAG methods. We also adopted the optimal hyperparameter configuration from LongRAG Zhao et al. (2024), and used identical hyperparameters across all methods: $k_{1}=100$ chunks retrieved initially, $k_{2}=7$ chunks selected after reranking, except in ablation studies. We maintain identical prompt templates across all methods to eliminate prompt engineering as a confounding variable.
For MacRAG-specific parameters, we varied the scale-up factor $\alpha\in\{1,4\}$ and the neighbor hop count $h\in\{0,1\}$, except during ablation studies. The ablation study and performance analysis in Section 4.2 demonstrate both robustness and the soundness of the architecture design of MacRAG, enhancing multi-hop reasoning capability.
This carefully controlled experimental framework ensures that observed performance differences can be attributed specifically to MacRAG’s architectural innovations, particularly hierarchical document representation and adaptive, multi-scale retrieval rather than variations in hyperparameter optimization or underlying retrieval technologies. By maintaining consistency while isolating architectural factors, we can precisely measure the impact of multi-scale retrieval on multi-hop reasoning performance.
4.2 Main Results
Model $\mathbf{k_{2}}$ Reranker R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_E&F Average
HotpotQA
LongRAG 7 bge-m3 67.67 67.99 68.96 64.30 68.49 66.73 66.10 67.18
MacRAG 7 bge-m3 67.59  (-0.08$\downarrow$) 68.63  (+0.64$\uparrow$) 70.53  (+1.57$\uparrow$) 65.56  (+1.26$\uparrow$) 70.29  (+1.80$\uparrow$) 66.59  (-0.14$\downarrow$) 67.32  (+1.22$\uparrow$) 68.07  (+0.89$\uparrow$, +1.32$\uparrow$)
LongRAG 12 bge-m3 68.57 67.65 70.31 64.74 70.14 67.63 67.49 68.08
MacRAG 12 bge-m3 67.88  (-0.69$\downarrow$) 69.87  (+2.22$\uparrow$) 70.72  (+0.41$\uparrow$) 65.98  (+1.24$\uparrow$) 70.66  (+0.52$\uparrow$) 69.05  (+1.42$\uparrow$) 67.84  (+0.35$\uparrow$) 68.86  (+0.78$\uparrow$, +1.15$\uparrow$)
2WikimultihopQA
LongRAG 7 bge-m3 59.36 65.56 68.27 55.31 67.36 64.42 63.88 64.45
MacRAG 7 bge-m3 62.32  (+2.96$\uparrow$) 66.34  (+0.78$\uparrow$) 71.63  (+3.36$\uparrow$) 69.72  (+14.41$\uparrow$) 73.98  (+6.62$\uparrow$) 66.61  (+2.19$\uparrow$) 72.63  (+8.75$\uparrow$) 69.03  (+4.58$\uparrow$, +7.11$\uparrow$)
LongRAG 12 bge-m3 60.08 66.77 69.28 58.50 69.39 64.90 65.28 64.89
MacRAG 12 bge-m3 64.29  (+4.21$\uparrow$) 67.20  (+0.43$\uparrow$) 70.95  (+1.67$\uparrow$) 69.46  (+10.96$\uparrow$) 73.90  (+4.51$\uparrow$) 67.49  (+2.59$\uparrow$) 71.80  (+6.52$\uparrow$) 69.30  (+4.41$\uparrow$, +6.80$\uparrow$)
Musique
LongRAG 7 bge-m3 42.34 48.08 47.88 42.17 47.92 43.70 44.06 45.16
MacRAG 7 bge-m3 45.54  (+3.20$\uparrow$) 46.68  (-1.40$\downarrow$) 49.58  (+1.70$\uparrow$) 46.76  (+4.59$\uparrow$) 49.53  (+1.61$\uparrow$) 47.94  (+4.24$\uparrow$) 49.14  (+5.08$\uparrow$) 47.88  (+2.72$\uparrow$, +6.02$\uparrow$)
LongRAG 12 bge-m3 41.53 46.96 49.58 41.60 49.57 46.76 46.33 46.05
MacRAG 12 bge-m3 46.44  (+4.91$\uparrow$) 45.81  (-1.15$\downarrow$) 51.02  (+1.44$\uparrow$) 46.70  (+5.10$\uparrow$) 51.54  (+1.97$\uparrow$) 50.25  (+3.49$\uparrow$) 49.50  (+3.17$\uparrow$) 48.75  (+2.70$\uparrow$, +5.86$\uparrow$)
Table 4: Robust test of MacRAG with various $k_{2}$ an alternative re-ranker on three multi-hop QA datasets, using GPT-4o and F1-score. The experiments were conducted with the same hyperparameters as reported in LongRAG Zhao et al. (2024).
($k_{2}$=7, marco-miniLM) R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_E&F Average
HotpotQA
LongRAG 65.46 69.40 66.60 63.32 66.20 66.08 66.09 66.16
MacRAG 67.15 (+1.69$\uparrow$) 69.55 (+0.15$\uparrow$) 69.00 (+2.40$\uparrow$) 65.44 (+2.12$\uparrow$) 68.52 (+2.32$\uparrow$) 66.36 (+0.28$\uparrow$) 66.53 (+0.44$\uparrow$) 67.51 (+1.35$\uparrow$, +2.04$\uparrow$)
- Propagation&Merging 64.17 (-1.29$\downarrow$) 68.91 (-0.49$\downarrow$) 67.82 (+1.22$\uparrow$) 68.06 (+4.74$\uparrow$) 68.44 (+2.24$\uparrow$) 63.77 (-2.31$\downarrow$) 65.57 (-0.52$\downarrow$) 66.68 (+0.52$\uparrow$, +0.79$\uparrow$)
- Scaling up 64.12 (-1.34$\downarrow$) 67.57 (-1.83$\downarrow$) 67.33 (+0.73$\uparrow$) 64.41 (+1.09$\uparrow$) 67.26 (+1.06$\uparrow$) 64.22 (-1.86$\downarrow$) 63.84 (-2.44$\downarrow$) 65.54 (-0.62$\downarrow$, -0.94$\downarrow$)
2WikimultihopQA
LongRAG 59.97 62.37 65.35 60.68 65.89 62.08 62.47 62.69
MacRAG 59.00 (-0.97$\downarrow$) 64.87 (+2.50$\uparrow$) 68.97 (+3.62$\uparrow$) 68.20 (+7.52$\uparrow$) 73.19 (+7.30$\uparrow$) 66.50 (+4.42$\uparrow$) 71.40 (+8.93$\uparrow$) 67.45 (+4.76$\uparrow$, +7.59$\uparrow$)
- Propagation&Merging 59.67 (-0.30$\downarrow$) 65.06 (+2.69$\uparrow$) 68.48 (+3.13$\uparrow$) 67.57 (+6.89$\uparrow$) 72.20 (+6.31$\uparrow$) 65.13 (+3.05$\uparrow$) 72.43 (+9.96$\uparrow$) 67.22 (+4.53$\uparrow$, +7.23$\uparrow$)
- Scaling up 59.00 (-0.97$\downarrow$) 61.89 (-0.48$\downarrow$) 67.63 (+2.28$\uparrow$) 60.41 (-0.27$\downarrow$) 67.03 (+1.14$\uparrow$) 66.37 (+4.29$\uparrow$) 65.77 (+3.30$\uparrow$) 64.01 (+1.32$\uparrow$, +2.11$\uparrow$)
Musique
LongRAG 38.98 41.90 43.64 37.72 43.83 42.00 41.97 41.43
MacRAG 44.76 (+5.78$\uparrow$) 45.74 (+3.84$\uparrow$) 47.42 (+3.78$\uparrow$) 47.80 (+10.08$\uparrow$) 50.09 (+6.26$\uparrow$) 48.57 (+6.57$\uparrow$) 51.00 (+9.03$\uparrow$) 47.91 (+6.48$\uparrow$, +15.63$\uparrow$)
- Propagation&Merging 41.61 (+2.63$\uparrow$) 45.43 (+3.53$\uparrow$) 48.00 (+4.36$\uparrow$) 47.48 (+9.76$\uparrow$) 52.26 (+8.43$\uparrow$) 40.09 (-1.91$\downarrow$) 45.84 (+3.87$\uparrow$) 45.82 (+4.39$\uparrow$, +10.58$\uparrow$)
- Scaling up 41.65 (+2.67$\uparrow$) 45.87 (+3.97$\uparrow$) 46.80 (+3.16$\uparrow$) 41.83 (+4.11$\uparrow$) 46.59 (+2.76$\uparrow$) 39.40 (-2.60$\downarrow$) 39.63 (-2.34$\downarrow$) 43.11 (+1.68$\uparrow$, +4.04$\uparrow$)
Table 5: Ablation Study of MacRAG with GPT-4o and F1-score. The experiments conducted with the same $k_{1}$, $k_{2}$ and reranker.
Figure 3: Performance trends across datasets for scale factors ($\alpha$).
Comparison with RAPTOR:  While RAPTOR Sarthi et al. (2024), a summarization and hierarchical clustering-based approach, shows gains on standard datasets, it struggles with retrieval quality on LongBench versions Bai et al. (2023) as shown in Table 2. Specifically, RAPTOR with Llama-3.1-8B shows performance decreases in 2WikimultihopQA and Musique, potentially associated with retrieval difficulty on Musique-Normal in (Yue et al., 2024). These results highlight the importance of both precise retrieval and effective long context construction for achieving consistent gains across various multi-hop reasoning and long-context datasets.
Integration with Multi-step Generation Method:  Integrating MacRAG with the LongRAG (E&F) generation method across multiple LLMs (Llama-3.1-8B, Gemini-1.5-pro, and GPT-4o) yields substantial performance improvements, as shown in Table 2. This underscores MacRAG’s versatility and robustness as a retrieval approach when paired with the LongRAG generation method. By constructing an effective long context, MacRAG demonstrates advantages in multi-step generation with LongRAG (E&F), producing consistent and meaningful gains across three datasets and three LLMs. In contrast, it is not straightforward to extend RAPTOR in combination with LongRAG (E&F). As further illustrated in Table 2, the performance gains for Gemini-1.5-pro and GPT-4o indicate that pairing MacRAG with more powerful LLMs can achieve even greater improvements, despite already high F1-scores obtained by other frameworks under these settings. We further observe that summarization during indexing improves retrieval quality. Our preliminary experiments show a 7.2% gain in precision when using summaries instead of raw slices.
Consistent Gains across Generation Methods:  In addition to MacRAG’s robust gains in both vanilla single-step (R&B) and multi-step generation (E&F), MacRAG consistently shows strong advantages across the four metrics listed in Table 3, covering all seven generation modes in Table 1. These substantial improvements across various datasets, metrics, and LLMs demonstrate the effectiveness of MacRAG’s long-context retrieval in enhancing both precision and recall for RAG. Notably, MacRAG achieves significant gains on the Musique datasets, which Yue et al. (2024) identifies as having challenging retrieval conditions. Beyond the average gains in these four metrics, Figure 2 further confirms MacRAG’s robust benefits in all seven single-/multi-step generation schemes in Table 1. Finally, regardless of the choice of $k_{2}$ or reranker, MacRAG maintains substantial advantages in all test settings in Table 4.
Enhanced Gains with Stronger Models: When applied to the more powerful Gemini-1.5-pro and GPT-4o models, MacRAG showcases its capacity to further improve performance by optimizing the retrieval process and building more relevant, informative contexts. Its use of hierarchical chunking and slicing, coupled with adaptive propagation and merging, maintains high precision while expanding coverage. This leads to more efficient handling of long documents, reduced computational overhead, and higher-quality generated answers overall.
Ablation Study:  Table 5 presents an ablation study demonstrating the effectiveness of MacRAG’s key components: Propagation and Merging, Scaling-up, and Hierarchical Slicing Retrieval. Each component contributes to cumulative performance gains. The combination of propagation and merging with the scaling-up mechanism increases the coverage of relevant contexts for multi-hop reasoning tasks. Figure 3 further illustrates the benefits of scaling up promising candidates with $\alpha=[2,3,4]$, showing how this approach effectively balances context lengths while incorporating additional relevant candidates. We observe that performance gains remain consistent across different $k_{2}$ values and rerankers, indicating robustness. Although the parameters $(k_{1},k_{2},\alpha,h)$ are fixed, MacRAG adaptively merges nearby slices depending on partial relevance signals. Removing this adaptivity leads to a 5% F1 drop, as shown in our ablations.
4.3 Generation Schemes and Input Lengths
Figure 4: Cumulative LLM’s input context lengths across steps for single/multi-step generation methods in Table 1.
While multi-step generation can improve answer quality, it also increases total input length across retrieval and generation steps for a single round. Figure 4 compares cumulative input lengths for single and multi-step generation methods in Table 1, given single retrieval with specific settings, highlighting that R&L, Full_Ext, and Full_E&F accumulate the most overhead by relying on entire documents. In contrast, MacRAG’s final $k_{2}$ chunks, though potentially larger than basic R&B chunks, preserve bounded context well below full-document with long lengths.
As Figure 2 and Table 5 illustrate, MacRAG consistently outperforms baselines under both R&B_Ext and R&B_E&F settings. Due to MacRAG’s advantage in effective and stable long context construction, even our intermediate R&B_E&F variant demonstrates significant improvements as shown in Table 5. While we do not report retrieval recall directly, MacRAG achieves competitive performance with reduced input length, suggesting effective evidence gathering through its hierarchical expansion.For instance, on 2WikimultihopQA, MacRAG (R&B_E&F) achieves an F1-score of 72.43 versus LongRAG (Full_E&F) at 65.89 (+6.54 absolute, +9.93% relative), using 8.19% less context. On Musique, the intermediate variant of MacRAG (R&B_E&F) reaches 45.84 compared to LongRAG (Full_E&F)’s 43.83 (+2.01 absolute, +4.58% relative) with a 45.87% reduction in context length.
4.4 Efficiency and Latency Analysis
MacRAG introduces a one-time offline summarization step during indexing like other methods (RAPTOR and SIRERAG). The summarization step for MacRAG provides both parallel standard RAG indexing, which does not affect real-time latency, and efficient maintenance.
Despite constructing richer multi-scale contexts, MacRAG maintains efficient retrieval and reranking, averaging 0.23 seconds across datasets, which is 38% faster than RAPTOR’s 0.37 seconds. The total inference time with Llama 3.1 8B is 0.99 seconds for MacRAG (0.23s retrieval + 0.76s generation), compared to 0.80 seconds for RAPTOR (0.37s + 0.43s) on typical multi-hop queries, balancing the added overhead with improved downstream accuracy on R&B.
During inference, MacRAG’s multi-scale adaptive expansions preserve the same level of computational complexity as baseline sub-second retrieval and reranking, thanks to shared reranking components and lightweight, index-based context construction.
Moreover, as shown in Table 5, Figure 4, and Subsection 4.3, MacRAG demonstrates improved efficiency over LongRAG in terms of the trade-off between final answer quality, input context length, and overall computational cost.
5 Conclusion
In this work, we introduce the Multi-scale Adaptive Context RAG (MacRAG) framework, which addresses the key trade-offs in RAG involving precision, coverage, and computational overhead for long-context question answering and multi-hop reasoning. By constructing query-adaptive, effective long contexts in real-time retrieval, MacRAG simultaneously improves precision and recall while maintaining efficiency. Experiments demonstrated its advantages with both single- and multi-step generation on multi-hop and large-document datasets, achieving notable performance gains. These findings underscore MacRAG’s potential as a core module for iterative and agentic RAG systems, offering substantial benefits for enterprise-level applications.
Acknowledgments
We are grateful to Prof. Xifeng Yan for his valuable feedback and suggestions throughout this research.
Appendix A Appendix
A.1 Adaptive Retrieval Mechanism Analysis
A key strength of MacRAG lies in its adaptive retrieval mechanism, which dynamically constructs query-specific contexts despite using fixed common hyperparameters ($k_{1}$, $k_{2}$) and MacRAG’s $\alpha$, and $h$. This adaptivity operates through several complementary mechanisms:
Query-dependent Adaptive Context Expansion:
For complex multi-hop queries, MacRAG naturally extends its context boundaries to encompass connecting information. For simpler queries, expansion remains more focused, selectively around the most promising content regions, but avoiding unnecessary content. Our ablation studies quantify the impact of this adaptivity. When the adaptivity of query-adaptive expansion of neighborhoods is removed by eliminating the propagation and merging steps (Table 5), performance drops by up to around 5% F1-score on datasets. This degradation is most pronounced on Musique, where connecting multiple scattered pieces of evidence is crucial. Table 5 additionally reveals another important aspect of query-adaptive expansion of borderline candidates "Scaling-up", and if we eliminate its corresponding "Scaling-up" component, then performance drops up to around 6% F1-score on datasets. This aligns with our expectation that adaptive expansion becomes increasingly valuable as query complexity grows.
A.2 Preprocessing Efficiency of MacRAG
Chunk Size and Token Budget:
All methods in our comparison use chunk sizes of approximately 400 tokens (around 1500 characters), which fall within the 200–500 token range used in prior work such as LongRAG. These sizes allow MacRAG to maintain sub-second retrieval and reranking times while keeping the context structure manageable for downstream language models.
Offline Summarization Cost:
MacRAG includes a one-time summarization step during indexing, which does not affect query-time latency. This step is comparable in cost to offline embedding computations in standard RAG pipelines. All runtime efficiency measurements reported here exclude this indexing cost, as it is amortized over the use of the document collection.
A.3 Latency Comparison between MacRAG and RAPTOR
Retrieval and Reranking Latency:
MacRAG maintains sub-second latency for retrieval and reranking across all evaluated datasets. Using dense retrieval with $k_{1}=100$ followed by cross-encoder reranking with $k_{2}=7$, the measured latency is 0.23 seconds on HotpotQA, 0.22 seconds on 2WikimultihopQA, and 0.24 seconds on Musique. In contrast, RAPTOR reports 0.43s, 0.24s, and 0.44s on the same datasets, averaging 0.37s. This represents a 38% speedup on average, even though both methods operate on the same top-100 retrieved chunks and use the same reranker (marco-miniLM). MacRAG achieves this by leveraging index-based merging after reranking, which reuses precomputed relevance scores and adds only a few milliseconds per query.
Generation Time and Trade-offs:
While retrieval and reranking are efficient, generation time varies depending on the context length produced by each method. MacRAG’s adaptive multi-scale expansion typically results in longer but more relevant contexts. With Llama 3.1 8B, the average generation time with R&B is 0.76 seconds, compared to 0.43 seconds for RAPTOR with R&B. This increase reflects MacRAG’s strategy of incorporating broader evidence to improve answer accuracy. Users seeking faster inference can reduce the expansion parameter $\alpha$ to trade off coverage for speed, though at the cost of some performance.
Total Latency Comparison:
Combining retrieval, reranking, and generation with R&B, the total end-to-end latency of MacRAG is 0.99 seconds (0.23s + 0.76s), compared to RAPTOR’s 0.80 seconds (0.37s + 0.43s). While MacRAG incurs a small additional cost, it consistently yields higher accuracy across multiple datasets and models, as shown in Table 2.
A.4 Efficiency Comparison between MacRAG and LongRAG
MacRAG and LongRAG both rely on index-based hierarchical retrieval and reranking over the same number of $k_{1}$ chunks. As a result, their retrieval and reranking steps show similar efficiency and latency. However, for the generation step, according to Table 5, Figure 4, and Subsection 4.3, MacRAG achieves improved efficiency compared to LongRAG in terms of the trade-off between final answer quality, input context length, and overall computational cost.
A.5 Discussion on Effectiveness of MacRAG’s Retrieval and Constructed Long Context
Enhanced Precision and Coverage:  MacRAG employs a hierarchical multi-scale strategy that begins with fine-grained slice retrieval and chunk-level re-ranking for precise identification of relevant content, then systematically incorporates broader document-level context through $h$-hop neighbor expansions. To preserve coherence without overwhelming the LLM, it strategically up-scales via the $(\alpha\times k_{2})$ factor, capturing borderline yet crucial segments. Rather than relying on compressed text or entire documents, MacRAG focuses on real-time construction of effective long contexts from promising original chunks while enforcing an upper bound on context length. By removing irrelevant portions from documents and maintaining a continuous, coherent subset of text, MacRAG avoids excessive token consumption, mitigates the “lost in the middle" phenomenon, and minimizes hallucinations, ultimately ensuring high recall for complex multi-hop queries.
A.6 Discussion on Single, Multi-Step, Iterative, and Agentic Generation with MacRAG
MacRAG’s modular design seamlessly supports single and multi-step generation, iterative retrieval, and agentic pipelines. By combining its core multi-scale retrieval with standard generation methods, MacRAG dynamically adapts context formation and refinement across multiple rounds without overwhelming the LLM. In single and multi-step modes (e.g., LongRAG Zhao et al. (2024) and standard multi-step QA in Section 3.3), MacRAG selects and assembles relevant text from large corpora while minimizing noise and retaining crucial connections, thereby improving both recall and precision for multi-hop reasoning. For iterative scenarios, such as IterDRAG Yue et al. (2024) or chain-of-RAG Wang et al. (2025), MacRAG can support efficient updates on the evidence set via $\mathcal{C}_{t}\leftarrow\mathrm{Merge}\Bigl{(}\mathcal{C}_{t-1},\,\mathrm{Retrieval}(q_{t})\Bigr{)}$ to unify old and newly retrieved content, promoting consistent coverage of bridging facts while discarding irrelevant material. In agentic RAG settings Asai et al. (2024); Jeong et al. (2024); Chen et al. (2025), MacRAG likewise prevents context explosion by focusing specifically on segments required for each action, thereby improving precision and recall over multiple steps. Our experiments (Section 4) confirm that integrating MacRAG with single and multi-step generation methods, potential core modules for multi-round advanced systems, consistently enhances complex multi-hop reasoning tasks and reduces error rates in long-context RAG.
A.7 Discussion on Graph-Enhanced MacRAG
A promising direction is extending MacRAG with a two-stage reranking strategy incorporating graph-based knowledge structures. After initial chunk-level reranking, MacRAG could perform a second reranking phase by expanding top candidates through both local index-based extensions and pre-constructed graph neighbors, enabling efficient coverage of both local and global relationships. This approach would leverage MacRAG’s bounded context guarantees at each phase while allowing controlled exploration of knowledge-guided connections, managing computational efficiency through bounded candidate sets during reranking. By applying MacRAG’s document-oriented indexing or relationship-based thresholding to merge expanded candidates, this strategy would enhance retrieval quality for complex queries requiring both precise local context and broad knowledge integration.
Figure 5: Performances of LongRAG and MacRAG regarding the fours metrics (Exact Match, F1-score, Precision, Recall) for three datasets (HotpotQA, 2WikimultihopQA, and Musique) and two LLMs (Gemini-1.5-pro and GPT-4o). Each row corresponds to a combination of dataset and LLM, and each column represents one of the metrics.
HotpotQA
Method ($k_{2}=7$) R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 63.59 63.93 61.93 63.38 61.95 57.86 57.53 61.45
MacRAG 63.02 (-0.57$\downarrow$) 66.76 (+2.83$\uparrow$) 64.90 (+2.97$\uparrow$) 65.96 (+2.58$\uparrow$) 64.40 (+2.45$\uparrow$) 60.59 (+2.73$\uparrow$) 62.14 (+4.61$\uparrow$) 63.97 (+2.52$\uparrow$, +4.10$\uparrow$)
2WikimultihopQA
Method ($k_{2}=7$) R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 60.13 62.99 57.82 58.31 58.17 53.12 53.54 57.73
MacRAG 58.38 (-1.75$\downarrow$) 63.48 (+0.49$\uparrow$) 60.90 (+3.08$\uparrow$) 66.74 (+8.43$\uparrow$) 64.75 (+6.58$\uparrow$) 55.32 (+2.20$\uparrow$) 64.19 (+10.65$\uparrow$) 61.85 (+4.12$\uparrow$, +7.14$\uparrow$)
Musique
Method ($k_{2}=7$) R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 34.90 40.97 33.73 35.01 34.04 29.82 28.96 33.92
MacRAG 43.31 (+8.41$\uparrow$) 43.41 (+2.44$\uparrow$) 40.68 (+6.96$\uparrow$) 45.81 (+10.80$\uparrow$) 43.34 (+9.30$\uparrow$) 40.61 (+10.79$\uparrow$) 42.76 (+13.80$\uparrow$) 42.84 (+8.92$\uparrow$), +26.30$\uparrow$)
Table 6: Experimental results of Gemini-1.5-pro with F1-score for comparing LongRAG vs MacRAG across HotpotQA, 2WikimultihopQA, and Musique datasets. The experiments conducted with the same hyper-parameter (k1=100 , k2=7) which is the reported best parameter of LongRAG Zhao et al. (2024). The columns represent various evaluation settings: R&B (Retrieval and Base), R&L (Retrieval and Long), Full_Ext (Extraction from Full Document), Fil (Filtering), Full_E&F (Extraction and Filtering combined), R&B_Ext (Extraction from Top-$k_{2}$ Chunks), and R&B_Ext_Fil (Extraction and Filtering from Top-$k_{2}$ Chunks). The absolute gains and relative percentage improvements from applying MacRAG to LongRAG are displayed in parentheses.
HotpotQA
$k_{2}=7$, marco-miniLM R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 65.46 69.40 66.60 63.32 66.20 66.08 66.09 66.16
MacRAG 67.15 (+1.69$\uparrow$) 69.55 (+0.15$\uparrow$) 69.00 (+2.40$\uparrow$) 65.44 (+2.12$\uparrow$) 68.52 (+2.32$\uparrow$) 66.36 (+0.28$\uparrow$) 66.53 (+0.44$\uparrow$) 67.22 (+1.06$\uparrow$)
$k_{2}=12$, marco-miniLM R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 67.97 69.80 67.91 63.56 67.82 66.91 66.44 67.20
MacRAG 68.41 (+0.44$\uparrow$) 69.64 (-0.16$\downarrow$) 69.95 (+2.04$\uparrow$) 65.42 (+1.86$\uparrow$) 69.09 (+1.27$\uparrow$) 67.57 (+0.66$\uparrow$) 66.95 (+0.51$\uparrow$) 68.00 (+0.80$\uparrow$)
$k_{2}=7$, bge-m3 R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 67.67 67.99 68.96 64.30 68.49 66.73 66.10 67.18
MacRAG 67.59 (-0.08$\downarrow$) 68.63 (+0.64$\uparrow$) 70.53 (+1.57$\uparrow$) 65.56 (+1.26$\uparrow$) 70.29 (+1.80$\uparrow$) 66.59 (-0.14$\downarrow$) 67.32 (+1.22$\uparrow$) 68.07 (+0.89$\uparrow$)
$k_{2}=12$, bge-m3 R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 68.57 67.65 70.31 64.74 70.14 67.63 67.49 68.08
MacRAG 67.88 (-0.69$\downarrow$) 69.87 (+2.22$\downarrow$) 70.72 (+0.41$\uparrow$) 65.98 (+1.24$\uparrow$) 70.66 (+0.52$\uparrow$) 69.05 (+1.42$\uparrow$) 67.84 (+0.35$\uparrow$) 68.86 (+0.78$\uparrow$)
2WikimultihopQA
$k_{2}=7$, marco-miniLM R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 59.97 62.37 65.35 60.68 65.89 62.08 62.47 62.69
MacRAG 59.00 (-0.97$\downarrow$) 64.87 (+2.50$\uparrow$) 68.97 (+3.62$\uparrow$) 68.20 (+7.52$\uparrow$) 73.19 (+7.30$\uparrow$) 66.50 (+4.42$\uparrow$) 71.40 (+8.93$\uparrow$) 67.45 (+4.76$\uparrow$)
$k_{2}=12$, marco-miniLM R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 62.64 65.43 70.67 62.06 70.66 67.56 66.60 66.52
MacRAG 63.47 (+0.83$\uparrow$) 67.06 (+1.63$\uparrow$) 70.61 (-0.06$\downarrow$) 67.84 (+5.78$\uparrow$) 72.24 (+1.58$\uparrow$) 67.83 (+0.27$\uparrow$) 72.23 (+5.63$\uparrow$) 68.75 (+2.23$\uparrow$)
$k_{2}=7$, bge-m3 R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 59.36 65.56 68.27 55.31 67.36 64.42 63.88 64.45
MacRAG 62.32 (+2.96$\uparrow$) 66.34 (+0.78$\uparrow$) 71.63 (+3.36$\uparrow$) 69.72 (+14.41$\uparrow$) 73.98 (+6.62$\uparrow$) 66.61 (+2.19$\uparrow$) 72.63 (+8.75$\uparrow$) 69.03 (+5.58$\uparrow$)
$k_{2}=12$, bge-m3 R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 60.08 66.77 69.28 58.50 69.39 64.90 65.28 64.89
MacRAG 64.29 (+4.21$\uparrow$) 67.20 (+0.43$\uparrow$) 70.95 (+1.67$\uparrow$) 69.46 (+10.96$\uparrow$) 73.90 (+4.51$\uparrow$) 67.49 (+2.59$\uparrow$) 71.80 (+6.52$\uparrow$) 69.30 (+4.41$\uparrow$)
Musique
$k_{2}=7$, marco-miniLM R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 38.98 41.90 43.64 37.72 43.83 42.00 41.97 41.43
MacRAG 44.76 (+5.78$\uparrow$) 45.74 (+3.84$\uparrow$) 47.42 (+3.78$\uparrow$) 47.80 (+10.08$\uparrow$) 50.09 (+6.26$\uparrow$) 48.57 (+6.57$\uparrow$) 51.00 (+9.03$\uparrow$) 47.77 (+6.34$\uparrow$)
$k_{2}=12$, marco-miniLM R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 41.20 43.85 48.19 40.48 47.22 44.66 44.51 44.30
MacRAG 45.85 (+4.65$\uparrow$) 47.66 (+3.81$\uparrow$) 48.27 (+0.08$\uparrow$) 49.01 (+8.53$\uparrow$) 48.99 (+1.77$\uparrow$) 49.35 (+4.69$\uparrow$) 50.57 (+6.06$\uparrow$) 48.53 (+4.23$\uparrow$)
$k_{2}=7$, bge-m3 R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 42.34 48.08 47.88 42.17 47.92 43.70 44.06 45.16
MacRAG 45.54 (+3.20$\uparrow$) 46.68 (-1.40$\downarrow$) 49.58 (+1.70$\uparrow$) 46.76 (+4.59$\uparrow$) 49.53 (+1.61$\uparrow$) 47.94 (+4.24$\uparrow$) 49.14 (+5.08$\uparrow$) 47.88 (+2.72$\uparrow$)
$k_{2}=12$, bge-m3 R&B R&L Ful_Ext Fil Full_E&F R&B_Ext R&B_Ext_Fil Average
LongRAG 41.53 46.96 49.58 41.60 49.57 46.76 46.33 46.05
MacRAG 46.44 (+4.91$\uparrow$) 45.81 (-1.15$\downarrow$) 51.02 (+1.44$\uparrow$) 46.70 (+5.10$\uparrow$) 51.54 (+1.97$\uparrow$) 50.25 (+3.49$\uparrow$) 49.50 (+3.17$\uparrow$) 48.75 (+2.70 $\uparrow$)
Table 7: Extensive experimental results for comparing LongRAG vs. MacRAG+LongRAG across HotpotQA, 2WikimultihopQA, and Musique datasets with two rerankers “marco-miniLM" and “bge-m3” via GPT-4o and F1-score. The experiments conducted with the same hyper-parameter ($k_{1}=100$, $k_{2}=7$) and ($k_{1}=100$, $k_{2}=12$) which is the reported best parameter of LongRAG Zhao et al. (2024). The columns represent various evaluation settings: R&B (Retrieval and Base), R&L (Retrieval and Long), Full_Ext (Extraction from Full Document), Fil (Filtering), Full_E&F (Extraction and Filtering combined), and R&B_Ext (Extraction from Top-$k_{2}$ Chunks). Gains from applying MacRAG to LongRAG (E&F) are displayed in parentheses with absolute gains on F1-scores and relative percentages of the improvements.

### v2

@@
 Abstract
-Long-context (LC) Large Language Models (LLMs) combined with Retrieval-Augmented Generation (RAG) hold strong potential for complex multi-hop and large-document tasks.
-However, existing RAG systems often suffer from imprecise retrieval, incomplete context coverage under constrained context windows, and fragmented information caused by suboptimal context construction.
-We introduce Multi-scale Adaptive Context RAG (MacRAG), a hierarchical retrieval framework that compresses and partitions documents into coarse-to-fine granularities, then adaptively merges relevant contexts through chunk- and document-level expansions in real time.
-By starting from the finest-level retrieval and progressively incorporating higher-level and broader context, MacRAG constructs effective query-specific long contexts, optimizing both precision and coverage.
-Evaluations on the challenging LongBench expansions of HotpotQA, 2WikiMultihopQA, and Musique confirm that MacRAG consistently surpasses baseline RAG pipelines on single- and multi-step generation with Llama-3.1-8B, Gemini-1.5-pro, and GPT-4o.
+Long-context large language models (LC LLMs) combined with retrieval-augmented generation (RAG) hold strong potential for complex multi-hop and large-document tasks.
+However, existing RAG systems often suffer from imprecise retrieval, incomplete context coverage under constrained windows, and fragmented information from suboptimal context construction.
+We introduce Multi-scale Adaptive Context RAG (MacRAG), a hierarchical RAG framework that compresses and partitions documents into coarse-to-fine granularities, then adaptively merges relevant contexts through real-time chunk- and document-level expansions.
+By initiating with finest-level retrieval and progressively incorporating broader, higher-level context, MacRAG constructs effective query-specific long contexts, optimizing both precision and coverage.
+Evaluations on challenging LongBench expansions of HotpotQA, 2WikiMultihopQA, and Musique confirm MacRAG consistently surpasses baseline RAG pipelines in single- and multi-step generation using Llama-3.1-8B, Gemini-1.5-pro, and GPT-4o.
 Our results establish MacRAG as an efficient, scalable solution for real-world long-context, multi-hop reasoning.
@@
 Figure 1: An overview of the MacRAG framework, consisting of two main phases: (1) top-down hierarchical indexing (upper), and (2) bottom-up multi-scale adaptive retrieval on the constructed hierarchy of document-chunk-summary-slice (lower).
-Large Language Models (LLMs) have substantially advanced question answering and reasoning, yet they frequently exhibit factual gaps or hallucinations when relying solely on internal parameters.
-Retrieval-Augmented Generation (RAG) (Guu et al., 2020; Lewis et al., 2020) mitigates these pitfalls by equipping LLMs with external evidence from large document collections.
-Meanwhile, Long Context (LC) LLMs, such as GPT-4o (OpenAI, 2024), Gemini 1.5 (Team et al., 2024), and Llama 3 (Dubey et al., 2024a), push context windows to hundreds of thousands or even millions of tokens, enabling an entire, particularly long context or document to be processed in principle.
-However, even these extended windows do not fully overcome the “Lost in the Middle” phenomenon (Liu et al., 2024; Xu et al., 2024b), wherein crucial mid-text information is overlooked, and their performance may degrade once length thresholds are surpassed (Leng et al., 2024; Yu et al., 2024).
-Within RAG, these issues manifest as three core trade-offs in document processing and retrieval for answer generation:
-  1. (1)
-    Context Length vs.
-Focus.
-Long contexts can improve recall but risk eclipsing critical details (particularly mid-text), whereas shorter contexts narrow in on precision but risk omitting essential evidence.
-  2. (2)
-    Chunking & Indexing. Finer-grained chunks support precise retrieval but can fragment continuity, while larger chunks preserve coherence but become redundant or unwieldy.
-  3. (3)
-    Coverage vs.
-Computation.
-Contexts across documents for multi-relational queries strain token budgets, latency limits or expensive pre-processing/updates, forcing a balance between broad coverage and efficiency on inference and system operations
-These trade-offs become especially acute in enterprise-grade or domain-specific deployments that demand “almost infinite” context for personalization or massive corpora.
-Moreover, recent iterative RAG approaches Yue et al.
-(2024); Wang et al.
-(2025) and agentic RAG pipelines Asai et al.
-(2024); Jeong et al.
-(2024); Chen et al.
-(2025) repeatedly invoke retrieval and generation, compounding inefficiencies unless the underlying retrieval module is robust.
-To systematically address these trade-offs, we propose Multi-scale Adaptive Context RAG (MacRAG), a hierarchical framework that unifies top-down offline indexing with bottom-up query-time retrieval.
-In an offline phase, MacRAG partitions each document into partially overlapping chunks, compresses them (e.g., via abstractive summarization), and further slices these compressed segments for fine-grained indexing via standard overlap chunking.
-This creates a hierarchical index from document level to fine-grained slices.
-At query time, MacRAG retrieves the most relevant slices at the finest level to ensure high precision, then progressively scales up to find more borderline contexts.
-It merges neighboring chunks and performs document-level expansions to construct effective long contexts in real time while controlling context size, addressing precision, coverage, and efficiency trade-offs simultaneously.
-This hierarchical multi-scale approach is compatible with both single- and multi-step generation methods (Xu et al., 2024a; Zhao et al., 2024) and integrates readily into iterative or agent-based RAG systems.
-Unlike prior works that improve either retrieval or ranking alone, MacRAG integrates structure-preserving indexing, bottom-up adaptive retrieval, and bounded context assembly for robust multi-hop reasoning.
-Through extensive evaluations on diverse long-document collections and multi-hop reasoning tasks, MacRAG demonstrates substantial improvements across various generation methods when used as a core RAG module with Llama-3.1-8B, Gemini-1.5-pro, and GPT-4o.
+Large language models (LLMs) have significantly advanced complex reasoning, but they still suffer from factual gaps or hallucinations when relying only on internal parameters (Zhao et al., 2024).
+Retrieval-Augmented Generation (RAG) addresses this by grounding LLMs in external evidence (Guu et al., 2020; Lewis et al., 2020).
+Long-context (LC) LLMs such as GPT-4o (OpenAI, 2024), Gemini 1.5 (Team et al., 2024), and Llama 3 (Dubey et al., 2024a) offer large input windows but remain limited.
+They often miss crucial mid-context information (Liu et al., 2024; Xu et al., 2024b) and their performance degrades at extreme context lengths (Leng et al., 2024; Yu et al., 2024).
+These issues in RAG systems give rise to three key trade-offs: (1) Context Length vs.
+Focus, where longer contexts improve recall but may obscure important details, while shorter ones enhance focus but risk omission of essential evidence; (2) Chunking and Indexing, where fine-grained chunks boost retrieval precision but harm coherence, while coarse chunks preserve structure but introduce redundancy; and (3) Coverage vs.
+Computation, where broader context improves reasoning but increases token cost and latency, especially in iterative or agentic pipelines (Yue et al., 2024; Asai et al., 2024).
+To address these interconnected trade-offs systematically, we propose Multi-scale Adaptive Context RAG (MacRAG).
+MacRAG integrates top-down offline indexing with bottom-up query-time adaptive retrieval.
+Offline, documents are partitioned into overlapping chunks, their content compressed via abstractive summarization, and these summaries are further sliced for fine-grained indexing.
+At query time, MacRAG retrieves precise slices, then adaptively reconstructs the context by merging these into parent chunks, incorporating neighboring chunks, and performing document-level expansions.
+This constructs effective, query-specific, and length-bounded long contexts, optimizing the balance between precision, coverage, and efficiency.
+MacRAG’s unified approach combines structure-preserving indexing, adaptive multi-scale retrieval, and bounded context assembly, offering a distinct, robust solution for complex multi-hop reasoning.
+Extensive experiments on challenging LongBench (Bai et al., 2023) datasets show significant gains over strong baselines using Llama-3.1-8B, Gemini-1.5-pro, and GPT-4o.
+This paper thus introduces a novel multi-scale retrieval architecture, empirically validates its significant benefits for demanding QA tasks, and presents an efficient, modular framework for advanced RAG applications.
 2 Related Work
-2.1 Multi-step Generation with Long Context in RAG
-Despite advances in LC LLMs, their integration with RAG systems continues to face the first trade-off discussed in Section 1.
-Recent research has explored context compression solutions to address the “Lost in the Middle" phenomenon and diminishing returns from excessive context.
-Recent approaches have explored two main techniques to effectively handle long contexts via multi-step generation with compression and filtering.
-Abstractive summarization methods, such as RECOMP-Abst Xu et al.
-(2024a), generate condensed text by rewriting retrieved passages.
-Extractive summarization approaches, including RECOMP-Extr Xu et al.
-(2024a), LLMLingua Jiang et al.
-(2023a, b), and EXIT Hwang et al.
-(2024), preserve original text through selective extraction with verification.
-LongRAG Zhao et al.
-(2024) offers a different perspective by exploring multiple strategies for utilizing retrieved chunks, comparing direct chunk usage against whole document incorporation and introducing a chain-of-thought-based filter and extractor for selecting relevant information.
-However, these post-retrieval multi-step generation approaches for long context face fundamental limitations.
-First, they rely exclusively on post-retrieval context compression without improving initial retrieval quality, creating an inherent performance ceiling.
-Second, generation based on compressed rather than original context risks information loss.
-Third, they incur significant computational overhead through multi-step processing on potentially extensive long context input, i.e., from corpora consisting of excessively long documents.
-In contrast, MacRAG focuses on the retrieval phase, efficiently constructing precise and comprehensive context inputs with bounded length for generation while minimizing related information loss.
-2.2 Hierarchical Retrieval Approaches for RAG
-Hierarchical retrieval approaches offer an alternative to post-retrieval compression by capturing longer or broader contexts during the retrieval phase.
-These methods organize information in hierarchical structures, aiming to improve both precision and recall.
-Sarthi et al.
-(2024) introduce RAPTOR, a tree-based architecture that recursively summarizes clustered text chunks for semantic matching.
-However, its exclusive focus on semantic similarity presents two key limitations: (1) it often misses important connected relationships that lack direct semantic similarity, hindering multi-hop reasoning capabilities, and (2) its rigid hierarchical structure can fragment interconnected knowledge.
-SIRERAG Zhang et al.
-(2025) extends this approach by incorporating relational connectivity alongside semantic similarity in a dual-tree framework.
-Although this enables more comprehensive knowledge synthesis, it introduces significant computational overhead and depends heavily on accurate proposition extraction and entity-based clustering.
-The system may miss subtle connections that fall outside its predefined relational model.
-Both approaches face substantial computational and practical limitations.
-Their clustering-based strategies incur high costs during database updates, which is particularly challenging for dynamic knowledge bases.
-Furthermore, these methods struggle with information overlap across hierarchical levels and lack robust mechanisms to balance semantic similarity with relational relevance.
-In contrast, MacRAG is more flexible and efficient by preserving the original content structure during indexing and retrieval.
-Recent work, such as GraphRAG (Edge et al., 2024) and HippoRAG (Gutiérrez et al., 2024), adopts hierarchical retrieval through symbolic graphs or sentence-level indexing, which improves coverage but introduces additional overhead.
-In contrast, MacRAG achieves similar benefits via dense, bottom-up expansion over pre-indexed multi-scale units without symbolic structure.
-Another line of work bypasses chunking altogether by retrieving entire passages or documents (Qian et al., 2024).
-While this simplifies retrieval, it reduces granularity control.
-MacRAG retains a flexible structure through hierarchical slicing and adaptive composition.
+Strategies for RAG with Long Context (LC) LLMs (OpenAI, 2024; Team et al., 2024; Dubey et al., 2024a) largely follow two directions.
+Post-retrieval context management aims to condense information after initial retrieval by employing methods such as abstractive summarization like RECOMP-Abst (Xu et al., 2024a) or extractive techniques including LLMLingua (Jiang et al., 2023) and RECOMP-Extr (Xu et al., 2024a).
+While these manage context size, they are bottlenecked by initial retrieval quality, risk information loss, and can be computationally intensive.
+MacRAG, by contrast, distinctly improves the retrieval phase itself, proactively constructing a high-quality, bounded context.
+Orthogonally, hierarchical retrieval approaches organize information during retrieval and ranking for broader coverage.
+Systems like GraphRAG (Edge et al., 2024) and HippoRAG (Gutiérrez et al., 2024) use symbolic graphs or sentence-level indexing, which can improve recall but often increase the overhead.
+RAPTOR (Sarthi et al., 2024) recursively summarizes clustered text chunks, potentially missing non-semantic relations, while SIRERAG (Zhang et al., 2025) further integrates relational connectivity at higher computational cost.
+LongRAG (Zhao et al., 2024) utilizes entire parent documents of top-ranked chunks alongside a multi-step generation scheme, a strategy that can be inefficient with very long document contexts.
+Based on Zhang et al.
+(2025)’s comparative analysis (Table 4 therein), which shows that GraphRAG underperforms on multi-hop QA while RAPTOR, HippoRAG, and SIRERAG attain similar F1-scores, we focus our evaluations on RAPTOR and LongRAG.
+Although RAPTOR’s semantic clustering may fragment knowledge and LongRAG’s full-document usage can be costly, both were selected for their promising trade-offs between strong performance and relatively lower overhead compared to other competitive methods, such as the noted underperformance of GraphRAG or the higher operational costs of SIRERAG.
+In contrast, MacRAG preserves original document structure through its offline hierarchical indexing and adaptively merges relevant and related contexts via its multi-scale retrieval and ranking at query time.
+This approach circumvents costly, repeated clustering and the need for explicit symbolic graphs, offering flexible assembly of multi-hop contexts with minimal overhead, and is designed with extendibility towards graph-based enhancements.
 3 Multi-scale Adaptive Context RAG
-We propose Multi-Scale Adaptive Context RAG (MacRAG), a hierarchical multi-scale adaptive retrieval system composed of three main components in sequence.
-First, the top-down hierarchical indexing processes documents via compression, chunking, and slicing, building a hierarchical index from coarse-grained documents to fine-grained slices.
-Second, the bottom-up multi-scale adaptive retrieval begins at the finest granularity and progressively traces back to broader contexts, merging relevant information to construct query-adaptive long contexts in real-time.
-Third, the response generation leverages these merged contexts to produce the final output.
-Our primary contribution lies in the retrieval pipeline (phases 1 and 2), which offers a modular foundation for integrating arbitrary response generation techniques in phase 3.
-Figure 1 concisely illustrates MacRAG, while Algorithm 1 describes its implementation, with mathematical details elaborated in the following sections.
-The retrieval process is adaptive in the sense that the amount and granularity of retrieved content vary depending on the query’s ambiguity and the initial slice-level relevance scores.
-3.1 Document Hierarchical Indexing
-Given a document corpus $\mathcal{D}=\{d_{1},\dots,d_{N}\}$, MacRAG performs offline processing to construct a hierarchy through chunking (§3.1.1), compression (§3.1.2), and slicing (§3.1.3), while building a slice-level vector DB at the finest granularity to enable retrieval.
-Instead of using raw sentences as the base unit, we use partial summaries to reduce redundancy and enhance semantic focus, which helps improve retrieval precision and reduces noise in the multi-scale expansion.
 Algorithm 1 MacRAG: Multi-scale Adaptive Context RAG
-Query $q$, document corpus $\mathcal{D}$, $k_{1}$ (for slice retrieval), $k_{2}$ (for final merged chunks), the number of hops $h$, up-scaling factor $\alpha$
+Query $q$, document corpus $\mathcal{D}$, $k_{1}$ (for slice retrieval), $k_{2}$ (for final merged chunks), the # of hops $h$, up-scaling factor $\alpha$
 Top-$k_{2}$ merged chunks $\mathcal{C}_{\mathrm{final}}$, generated answer
@@
 (1.4) Encode each slice and store its embedding & metadata (doc ID, chunk ID, offset, etc) in the database.
-Phase 2: Bottom-up Multi-Scale Adaptive Retrieval
+Phase 2: Bottom-up Multi-scale Adaptive Retrieval and Ranking
 (2.1) Retrieve the top-$k_{1}$ slices based on similarity to $q$.
@@
 Provide $\mathcal{C}_{\mathrm{final}}$ as context to an LLM (or any downstream model) to generate the final answer.
+We propose Multi-scale Adaptive Context RAG (MacRAG), a hierarchical multi-scale adaptive retrieval system with three sequential components.
+First, top-down hierarchical indexing compresses documents, partitions them into chunks, and slices them to build a hierarchical index from coarse-grained documents to fine-grained slices with multi-level indices.
+Second, bottom-up multi-scale adaptive retrieval starts from the finest granularity to ensure precision, and progressively expands to broader contexts by merging relevant information to construct query-specific long contexts dynamically in real time.
+Third, the response generation leverages these carefully assembled contexts to produce the answer.
+Our main contribution lies in the retrieval, ranking, and adaptive input context construction pipeline (phases 1 and 2), which serves as a modular foundation for integrating any response generation method in phase 3.
+Figure 1 provides an overview of MacRAG, and Algorithm 1 outlines the methods along with equations in subsequent sections.
+3.1 Document Hierarchical Indexing
+Given a document corpus $\mathcal{D}=\{d_{1},\dots,d_{N}\}$, MacRAG performs offline processing to construct a hierarchy through chunking (§3.1.1), compression (§3.1.2), and slicing (§3.1.3), while building a slice-level vector DB at the finest granularity to empower precise retrieval.
+This strategy aims to reduce redundancy and sharpen semantic focus, thereby enhancing retrieval precision and minimizing noise during the subsequent multi-scale expansion phase.
 3.1.1 Chunking
-Each document $d\in\mathcal{D}$ is split into partially overlapping chunks,
-$\mathcal{C}_{d}=\{c_{1},c_{2},\dots,c_{m}\},$ (1)
-where each chunk $c_{i}$ spans approximately 200 to 500 tokens, with a fixed overlap of 10–20 tokens between consecutive chunks.
+Each document $d\in\mathcal{D}$ is split into partially overlapping chunks, $\mathcal{C}_{d}=\{c_{1},c_{2},\dots,c_{m}\}$, where each chunk $c_{i}$ spans approximately 200 to 500 tokens, with a fixed overlap of 10–200 tokens between consecutive chunks.
 This controlled overlap preserves semantic continuity across chunk boundaries.
 Although the chunking is token-based by default, it can alternatively be configured at the character level.
+We applied the similar range of chunk sizes used by LongRAG  (Zhao et al., 2024), more details in Appendix A.5.
 3.1.2 Compression
-Each chunk $c_{i}$ is transformed into a more compressed summary $s_{i}$ and thus we obtain a set of compressed summaries $\mathcal{S}_{d}=\{\,s_{1},\dots,s_{m}\}$, where $s_{i}=\mathrm{Compress}(c_{i})$ and $\mathrm{Compress}(\cdot)$ represents a summarization model.
+Each chunk $c_{i}$ is transformed into a more compressed summary $s_{i}$, yielding a set of summaries $\mathcal{S}_{d}=\{s_{1},\dots,s_{m}\}$, where $s_{i}=\mathrm{Compress}(c_{i})$.
+The $\mathrm{Compress}(\cdot)$ represents a summarization model.
 It can be either extractive (selecting salient sentences) or abstractive (using a generative model to generate a summary).
 By default, we use an LLM to generate the abstractive summarization.
-This step reduces redundancy while retaining core factual and named-entity information.
+This step effectively reduces redundancy while retaining core factual information.
 3.1.3 Slicing
-To facilitate finer-grained retrieval, each summary $s_{i}$ is further split into overlapping slices: $\mathcal{L}_{i}=\bigl{\{}\,\ell_{i,1},\ell_{i,2},\dots,\ell_{i,j},\dots\bigr{\}}$, where each slice $\ell_{i,j}$ typically spans $50\text{-}200$ tokens with partial overlap as in the standard overlap chunking.
+To facilitate finer-grained retrieval, each summary $s_{i}$ is further split into overlapping slices: $\mathcal{L}_{i}=\bigl{\{}\,\ell_{i,1},\ell_{i,2},\dots,\ell_{i,j},\dots\bigr{\}}$, where each slice $\ell_{i,j}$ typically spans $50\text{-}200$ tokens with partial overlap as in the standard overlap chunking, and our parameter details are in Appendix A.5.
 Each slice is then embedded via $\mathbf{e}_{i,j}=\mathrm{Embed}(\ell_{i,j})$, where $\mathrm{Embed}(\cdot)$ represents a text encoder (e.g., multilingual-e5-large).
@@
 MacRAG can optionally index chunk and summary embeddings for multi-view retrieval.
-3.2 Multi-Scale Adaptive Retrieval
-Given an input query $q$, MacRAG implements a multi-scale adaptive retrieval pipeline over the constructed hierarchy through five sequential steps, ultimately selecting the top-$k_{2}$ merged chunks with contextually relevant neighbor expansions for enhanced retrieval precision and recall.
+3.2 Multi-scale Adaptive Retrieval and Ranking
+Given an input query $q$, MacRAG implements a multi-scale adaptive retrieval pipeline over the constructed hierarchy through five sequential steps, ultimately selecting the top-$k_{2}$ merged chunks with contextually relevant neighbor expansions for enhanced both precision and recall.
 3.2.1 Initial Slice-level Retrieval
-With slices serving as the finest granular units in our hierarchical system and vector DB, we first compute the query-slice relevance scores for each slice $\ell_{i,j}$ in the vector DB:
-$r_{q,\ell_{i,j}}=\mathrm{Rel}\!\bigl{(}q,\,\ell_{i,j}\bigr{)},$ (2)
-where $\mathrm{Rel}(\cdot,\cdot)$ can be implemented with relevance model, e.g., cosine similarity and $L_{2}$ distance in dense retrieval, sparse retrieval, hybrid retrieval, and so on.
+With slices serving as the finest granular units in our hierarchical system and vector DB, we first compute the query-slice relevance scores for each slice $\ell_{i,j}$ in the vector DB: $r_{q,\ell_{i,j}}=\mathrm{Rel}\!\bigl{(}q,\,\ell_{i,j}\bigr{)}$, where $\mathrm{Rel}(\cdot,\cdot)$ can be implemented with relevance model, e.g., cosine similarity and $L_{2}$ distance in dense retrieval, sparse retrieval, hybrid retrieval, and so on.
 We then retrieve the top-$k_{1}$ most relevant slices.
-This fine-grained retrieval ensures high precision by identifying slices with strong semantic alignment to query $q$.
+This fine-grained retrieval ensures high precision by identifying slices with strong semantic alignment to the query $q$.
 3.2.2 Parent Chunk Mapping
-Each retrieved slice corresponds to a parent chunk $c_{i}$. To prevent redundancy when multiple retrieved slices originate from the same chunk, we perform a unique mapping:
-$\mathcal{C}_{k_{1}}^{{}^{\prime}}=\mathrm{unique}\bigl{(}{c_{i}\mid\ell_{i,\cdot}\in\mathrm{top}\text{-}k_{1}(r_{q,\ell})}\bigr{)},$ (3)
-where we identify and consolidate the parent chunks of all retrieved slices and obtain a set of selected chunks used for the next steps.
+Each retrieved slice corresponds to a parent chunk $c_{i}$.
+To prevent redundancy when multiple retrieved slices originate from the same chunk, we perform a unique mapping: $\mathcal{C}_{k_{1}}^{{}^{\prime}}=\mathrm{unique}\bigl{(}{c_{i}\mid\ell_{i,\cdot}\in\mathrm{top}\text{-}k_{1}(r_{q,\ell})}\bigr{)}$, where we identify and consolidate the parent chunks of all retrieved slices and obtain a set of selected chunks used for the next steps.
 3.2.3 Chunk-level Re-Ranking
-For each identified unique relevant chunk from the mapping as $c_{i}\in\mathcal{C}_{k_{1}}^{{}^{\prime}}$ to the query, we compute a chunk-level ranking score for the query:
-$r_{q,c_{i}}=\mathrm{ReRank}\bigl{(}q,\,c_{i}\bigr{)},$ (4)
-utilizing a cross-encoder (e.g., marco-miniLM) or advanced ranking model that evaluates the complete chunk content, mitigating potential “information fragmentation" inherent in slice-level retrieval.
+For each identified unique relevant chunk from the mapping as $c_{i}\in\mathcal{C}_{k_{1}}^{{}^{\prime}}$ to the query, we compute a chunk-level ranking score for the query: $r_{q,c_{i}}=\mathrm{ReRank}\bigl{(}q,\,c_{i}\bigr{)}$, utilizing a cross-encoder (e.g., marco-miniLM) or advanced ranking model that evaluates the complete chunk content, mitigating potential “information fragmentation" inherent in slice-level retrieval.
 The chunks in $\mathcal{C}_{k_{1}}^{{}^{\prime}}$ are then reordered according to their ranking scores to the query $q$.
 3.2.4 Scaling-Up & Document-level Ranking
-In this step, we aim to select content segments, which can be used to construct $k_{2}$ precise long contexts with upper bounds in the later step to ensure both relevance and sufficient context coverage for final response generation with LLM.
-To achieve this, we incorporate top-ranked chunks and their document neighbors, thereby mitigating information fragmentation and enabling multi-hop reasoning.
-The process consists of the following steps.
-Scaled Top Chunk Selection:  Instead of picking exact top-$k_{2}$ chunks, we consider more borderline chunk candidates from the re-ranking stage, since the computed $r_{q,c_{i}}$ may not be perfect.
-Thus, MacRAG adopts scaling factor $\alpha$ and selects the top-($k_{2}\times\alpha$) chunks from $\mathcal{C}_{k_{1}}^{{}^{\prime}}$ for up-scaling with $r_{q,c_{i}}$ in Equation 4:
+This step aims to select optimal content segments for constructing $k_{2}$ precise, bounded long contexts for the final LLM generation, ensuring both relevance and sufficient coverage.
+It incorporates top-ranked chunks and their context neighbors in the same document to mitigate information fragmentation and support multi-hop reasoning.
 […]
-where $\alpha\geq 1$ and $(k_{2}\times\alpha)\leq|\mathcal{C}^{{}^{\prime}}_{k_{1}}|$.
-$\alpha=1$ does not scale-up the number of borderline candidates, but moderate higher $\alpha$ (e.g., 3 or 4) can provide higher probabilities to include borderline but crucial chunks to improve coverage while maintaining certain-level of qualities for candidates.
-To balance among precision, recall, and context lengths, $\alpha\in\{2,3,4\}$ would be promising where $(k_{2}\times\alpha)\leq|\mathcal{C}^{{}^{\prime}}_{k_{1}}|$, and we will also demonstrate that selection $\alpha$ is not sensitive, but robust with handling trade-off problems well in Section 4.
-This approach can also potentially facilitate bridging multi-hop queries.
-Document Selection:  We further track one level up from the $(k_{2}\times\alpha)$ chunks $c_{i}$ in $\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$ by mapping $c_{i}$ into their parents to identify unique candidates for selecting top-$k_{2}$ documents.
-We then compute the rank score $r_{q,d}$ of mapped document with re-ranking mechanism leveraging chunks’ rank score $r_{q,c_{i}}$ from all the chunks $c_{i}$ within the document $d$ and $c_{i}$ in $\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$:
-$r_{q,d}\;=\;\mathrm{ReRank_{Doc}}\Bigl{(}\{c_{i}\in d~{}and~{}c_{i}\in\mathcal{C}^{{}^{\prime}}_{q,k_{2}}\}\Bigr{)}.$ (6)
-Potential choices for document ranking functions could be mean, max, sum, or an LLM-based ranker.
-If chunk’s rank score $r_{q,c_{i}}$ is computed by using a cross encoder, then it can be either positive and negative, then $\max({r_{q,c_{i}}})$ would be the base function for document ranking.
-Based on the computed $r_{q,d}$, we construct the $\mathcal{D}_{q,k_{2}}$ consisting of top-$k_{2}$ documents.
-We ensure top-$k_{2}$ distinct docs so coverage is not restricted to a single or few docs.
-If we consider properly higher $\alpha$ (e.g., 3 or 4 depending on the overlapped ratio of original chunks), then there are fewer risks for unfilled top-ranked documents.
-This is critical when relevant evidence is scattered across multiple documents.
-3.2.5 Neighbor Chunk Propagation & Merging
+Document Selection.
+From the scaled set of $(k_{2}\times\alpha)$ chunks in $\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$, we identify their unique source documents.
+To select the top-$k_{2}$ documents, we compute a document-level rank score $r_{q,d}$ for each candidate document $d$.
+This score $r_{q,d}$ leverages the chunk-level scores $r_{q,c_{i}}$ of document’s all constituent chunks $c_{i}$ that are also present in $\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$: $r_{q,d}=\mathrm{ReRank_{Doc}}\Bigl{(}\{c_{i}\in d\text{~{}and~{}}c_{i}\in\mathcal{C}^{{}^{\prime}}_{q,k_{2}}\}\Bigr{)}$.
+Document ranking functions can include mean, max, sum of chunk scores, or an LLM-based ranker (with $\max(r_{q,c_{i}})$ being a suitable base if using cross-encoder scores, which can be either positive or negative).
+Based on $r_{q,d}$, we form $\mathcal{D}_{q,k_{2}}$, a set of top-$k_{2}$ distinct documents.
+Ensuring document diversity, especially with appropriate $\alpha$ values (e.g., 3 or 4 depending on the overlapped ratio of original chunks), mitigates the risk of concentrating on too few sources and is critical when evidence is scattered across multiple documents.
+3.2.5 Neighbor Chunk Propagation & Merge for Input Context Construction
 Within each of the top-$k_{2}$ distinct documents, we focus on chunks $c_{i}$ both in $\mathcal{D}_{q,k_{2}}$ and $\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$, as these represent the most relevant content segments.
-For such a chunk $c_{i}$, we define $\mathcal{N}_{h}(c_{i})$ which consists of $c_{i}$’s neighbors within $h$-hops and $c_{i}$ itself, and process extension of $c_{i}$ to get $\mathcal{N}(c_{i})$ by using its associated indices,
-$\mathcal{N}_{h}(c_{i})=\;\bigl{\{}\,c_{i-h},\dots,c_{i},\dots,c_{i+h}\bigr{\}}.$ (7)
-We then merge $\mathcal{N}_{h}(c_{i})$ within the same ranked document to form the final merged chunk for adaptive long context to the query $q$:
-$c^{merge}_{q,d}~{}=~{}$ $\mathrm{Merge}\bigl{(}\mathcal{N}_{h}(c_{i})\bigr{)},$ (8)
-$\text{where}~{}d\in$ $\mathcal{D}_{q,k_{2}},~{}c_{i}\in d~{}and~{}c_{i}\in\mathcal{C}^{{}^{\prime}}_{q,k_{2}}.$
-As described in Equation 8, from the top-$k_{2}$ documents in $\mathcal{D}_{q,k_{2}}$, we can select top-$k_{2}$ merged chunks $c^{merge}_{q,d}$ by considering its order along with top document order with $r_{q,d}$.
-Then, finally we have top-$k_{2}$ merged chunks as
-$\mathcal{C}_{\mathrm{final}}~{}=~{}\mathrm{TopK}_{k_{2}}\!\Bigl{(}c^{merge}_{q,d}\Bigr{)}.$ (9)
+For such a chunk $c_{i}$, we define $\mathcal{N}_{h}(c_{i})$ which consists of $c_{i}$’s neighbors within $h$-hops and $c_{i}$ itself, and process extension of $c_{i}$ to get $\mathcal{N}(c_{i})$ by using its associated indices, $\mathcal{N}_{h}(c_{i})=\;\bigl{\{}\,c_{i-h},\dots,c_{i},\dots,c_{i+h}\bigr{\}}$.
+We then merge $\mathcal{N}_{h}(c_{i})$ within the same ranked document to form the final merged chunk for adaptive long context to the query $q$: $c^{merge}_{q,d}=\mathrm{Merge}\bigl{(}\mathcal{N}_{h}(c_{i})\bigr{)}$, where $d\in\mathcal{D}_{q,k_{2}}$, $c_{i}\in d$, and $c_{i}\in\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$.
+From the top-$k_{2}$ documents in $\mathcal{D}_{q,k_{2}}$, we can select top-$k_{2}$ merged chunks $c^{merge}_{q,d}$ by considering their order along with the top document order with $r_{q,d}$.
+Then, finally, we have top-$k_{2}$ merged chunks as $\mathcal{C}_{\mathrm{final}}=\mathrm{TopK}_{k_{2}}\!\Bigl{(}c^{merge}_{q,d}\Bigr{)}$.
 The final merged chunks $\mathcal{C}_{\text{final}}$ provide diverse views across the top-$k_{2}$ documents while optimizing context length (Zhao et al., 2024), combining the most relevant content with minimal redundancy.
-Mode Description
-R&B Single-step generation using retrieved top-$k_{2}$ chunks, which is the vanilla basic version.
-R&L Single-step generation using complete documents from top-$k_{2}$ chunks. Prone to “Lost in the Middle” effect and inefficient for long documents.
-Full_Ext Multi-step generation: (1) LLM extracts query-relevant content from full documents of top-$k_{2}$ chunks, (2) Generates answer using both chunks and extracted content.
-Fil Filtering re-ranked top-$k_{2}$ chunks, e.g., via LLM. Improves precision but may lose borderline info.
-Full_E&F Multi-step generation combining Full_Ext’s extracted content and Fil’s filtered chunks. Balances coverage and precision, but expensive for long doc.
-R&B_Ext Multi-step generation with content extraction limited to top-$k_{2}$ chunks, reducing computational overhead.
-R&B_E&F Multi-step generation combining R&B_Ext’s extracted content and Fil’s filtered chunks. Optimizes precision-recall trade-off w/o processing full doc.
-Table 1: Seven different generation methods with single-step and multi-step generation schemes using retrieved top-$k_{2}$ chunks.
 3.3 Single/Multi-step Post Retrieval Generation
 Our retrieval framework is modular, enabling integration with various post-retrieval generation approaches.
-These include both vanilla single-step generation directly using the retrieved top-$k_{2}$ chunks in a single pass and iterative multi-step generation as in LongRAG (Zhao et al., 2024), which introduces CoT-guided filter (which checks chunk-level relevance) and LLM-augmented extractor (which locates global context in original passages).
+These include both vanilla single-step generation directly using the retrieved top-$k_{2}$ chunks in a single pass and iterative multi-step generation as in LongRAG (Zhao et al., 2024), which introduces a CoT-guided filter (which checks chunk-level relevance) and an LLM-augmented extractor (which locates global context in original passages).
 Following the comprehensive exploration of single and multi-step generation variants in LongRAG, we evaluate MacRAG’s performance when combined with different post-retrieval generation methods.
 Specifically, we investigate seven generation methods, as detailed in Table 4.
-Five of these methods—R&B, R&L, Full_Ext, Fil, and Full_E&F—are adopted directly from LongRAG.
+Five of these methods (R&B, R&L, Full_Ext, Fil, and Full_E&F) are adopted directly from LongRAG.
 Among these, R&L, Full_Ext, and Full_E&F process the entire set of top-$k_{2}$ documents, which can introduce computational overhead when handling very long texts.
@@
 These processes are performed once during offline preparation, eliminating query-time overhead.
-Moreover, this document-level independence simplifies updates and streamlines database operations, avoiding the complexities associated with hierarchical multi-clustering Sarthi et al.
+Moreover, this document-level independence simplifies updates and streamlines database operations, avoiding the complexity associated with hierarchical multi-clustering Sarthi et al.
 (2024); Zhang et al.
@@
 A bounded context size is enforced through parameters $(k_{2}\times\alpha)$ and chunk dimensions, ensuring scalability to enterprise-level corpora without degrading performance.
-This controlled scalability, along with real-time context assembly, makes MacRAG highly valuable for massive long-document collections, “almost infinite context” in advanced personalization, and iterative and agentic RAGs.
+This controlled scalability and real-time context assembly underscore MacRAG’s value for massive long-document collections, “almost infinite context" scenarios in advanced personalization, and as a robust foundation for iterative RAG or agentic systems.
 Model HotpotQA 2WikimultihopQA Musique Average
@@
 Recall 35.69 45.16 (+9.46$\uparrow$, +26.50$\uparrow$) 45.06 51.67 (+6.61$\uparrow$, +14.67$\uparrow$)
-Table 3: Consolidated results for four performance metrics across three datasets, comparing LongRAG and MacRAG using the Gemini-1.5-pro and GPT-4o models.
+Table 2: Consolidated results for four metrics across three datasets, comparing LongRAG and MacRAG using the Gemini-1.5-pro and GPT-4o models.
 The results are averaged over seven generation settings in Table 4.
@@
 4.1 Experimental Setup
-Benchmark Datasets and Challenges:
-We evaluate MacRAG on three multi-hop question answering benchmarks from the expanded LongBench framework (Bai et al., 2023): HotpotQA, 2WikimultihopQA, and Musique.
-These datasets have become the standard evaluation suite for hierarchical retrieval research, including recent systems such as SIRERAG Zhang et al.
+We evaluate MacRAG on three challenging multi-hop question answering benchmarks from LongBench (Bai et al., 2023): HotpotQA, 2WikiMultihopQA, and Musique.
+These datasets, with their tangential passages and obscured connecting information, specifically test a system’s ability to handle the "Lost in the Middle" phenomenon and perform robust multi-hop reasoning—precisely the challenges MacRAG addresses.
+Section A.2 in the Appendix provides a detailed explanation about these datasets.
+Models and Evaluations.
+Our evaluations employ three prominent LC LLMs: GPT-4o (OpenAI, 2024), Gemini-1.5-pro (Team et al., 2024), and the open-source Llama-3.1-8B-instruct (Dubey et al., 2024a).
+Performance is measured using Exact Match (EM), F1-score, Precision, and Recall.
+We primarily compare MacRAG against strong baselines RAPTOR (Sarthi et al., 2024) and LongRAG (Zhao et al., 2024), selected for their representative state-of-the-art performance and efficiency balance in hierarchical RAG (as discussed in Section 2).
+To ensure rigorous comparative analysis across all RAG methods (MacRAG, LongRAG, and RAPTOR) use the same reranker (marco-miniLM), and LongRAG’s reported optimal hyperparameters ($k_{1}=100$ initial chunks, $k_{2}=7$ final chunks, unless in ablation).
+We also applied seven different generation methods described in Appendix A.1, and Table 4 for detailed comparisons.
+MacRAG-specific parameters are scale-up factor $\alpha\in\{1,4\}$ and the neighbor hop count $h\in\{0,1\}$, except during ablation studies.
+The ablation study and performance analysis in Section 4.2 demonstrate both robustness and the soundness of the architecture design of MacRAG, enhancing multi-hop reasoning capability.
+This controlled setup ensures that performance differences result from MacRAG’s architectural innovations, particularly its hierarchical representation and adaptive multi-scale retrieval, rather than from variations in hyperparameters or retrieval components.
+By isolating architectural factors, we can clearly assess the impact of multi-scale retrieval on multi-hop reasoning.
+4.2 Main Results
+($k_{2}$=7, marco-miniLM) R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_E&F Average
+HotpotQA
+LongRAG 65.46 69.40 66.60 63.32 66.20 66.08 66.09 66.16
+MacRAG 67.15 (+1.69$\uparrow$) 69.55 (+0.15$\uparrow$) 69.00 (+2.40$\uparrow$) 65.44 (+2.12$\uparrow$) 68.52 (+2.32$\uparrow$) 66.36 (+0.28$\uparrow$) 66.53 (+0.44$\uparrow$) 67.51 (+1.35$\uparrow$, +2.04$\uparrow$)
+- Propagation&Merging 64.17 (-1.29$\downarrow$) 68.91 (-0.49$\downarrow$) 67.82 (+1.22$\uparrow$) 68.06 (+4.74$\uparrow$) 68.44 (+2.24$\uparrow$) 63.77 (-2.31$\downarrow$) 65.57 (-0.52$\downarrow$) 66.68 (+0.52$\uparrow$, +0.79$\uparrow$)
+- Scaling up 64.12 (-1.34$\downarrow$) 67.57 (-1.83$\downarrow$) 67.33 (+0.73$\uparrow$) 64.41 (+1.09$\uparrow$) 67.26 (+1.06$\uparrow$) 64.22 (-1.86$\downarrow$) 63.84 (-2.44$\downarrow$) 65.54 (-0.62$\downarrow$, -0.94$\downarrow$)
+2WikimultihopQA
+LongRAG 59.97 62.37 65.35 60.68 65.89 62.08 62.47 62.69
+MacRAG 59.00 (-0.97$\downarrow$) 64.87 (+2.50$\uparrow$) 68.97 (+3.62$\uparrow$) 68.20 (+7.52$\uparrow$) 73.19 (+7.30$\uparrow$) 66.50 (+4.42$\uparrow$) 71.40 (+8.93$\uparrow$) 67.45 (+4.76$\uparrow$, +7.59$\uparrow$)
+- Propagation&Merging 59.67 (-0.30$\downarrow$) 65.06 (+2.69$\uparrow$) 68.48 (+3.13$\uparrow$) 67.57 (+6.89$\uparrow$) 72.20 (+6.31$\uparrow$) 65.13 (+3.05$\uparrow$) 72.43 (+9.96$\uparrow$) 67.22 (+4.53$\uparrow$, +7.23$\uparrow$)
+- Scaling up 59.00 (-0.97$\downarrow$) 61.89 (-0.48$\downarrow$) 67.63 (+2.28$\uparrow$) 60.41 (-0.27$\downarrow$) 67.03 (+1.14$\uparrow$) 66.37 (+4.29$\uparrow$) 65.77 (+3.30$\uparrow$) 64.01 (+1.32$\uparrow$, +2.11$\uparrow$)
+Musique
+LongRAG 38.98 41.90 43.64 37.72 43.83 42.00 41.97 41.43
+MacRAG 44.76 (+5.78$\uparrow$) 45.74 (+3.84$\uparrow$) 47.42 (+3.78$\uparrow$) 47.80 (+10.08$\uparrow$) 50.09 (+6.26$\uparrow$) 48.57 (+6.57$\uparrow$) 51.00 (+9.03$\uparrow$) 47.91 (+6.48$\uparrow$, +15.63$\uparrow$)
+- Propagation&Merging 41.61 (+2.63$\uparrow$) 45.43 (+3.53$\uparrow$) 48.00 (+4.36$\uparrow$) 47.48 (+9.76$\uparrow$) 52.26 (+8.43$\uparrow$) 40.09 (-1.91$\downarrow$) 45.84 (+3.87$\uparrow$) 45.82 (+4.39$\uparrow$, +10.58$\uparrow$)
+- Scaling up 41.65 (+2.67$\uparrow$) 45.87 (+3.97$\uparrow$) 46.80 (+3.16$\uparrow$) 41.83 (+4.11$\uparrow$) 46.59 (+2.76$\uparrow$) 39.40 (-2.60$\downarrow$) 39.63 (-2.34$\downarrow$) 43.11 (+1.68$\uparrow$, +4.04$\uparrow$)
+Table 3: Ablation study of MacRAG with GPT-4o and F1-score, using the same $k_{1}$, $k_{2}$, and reranker.
+Figure 3: Performance trends across datasets for scale factors ($\alpha$).
+Comparison with RAPTOR.
+While RAPTOR Sarthi et al.
+(2024), a summarization and hierarchical clustering-based approach, shows gains on standard datasets, it struggles with retrieval quality on LongBench versions Bai et al.
+(2023) as shown in Table 1.
+Specifically, RAPTOR with Llama-3.1-8B shows performance decreases in 2WikimultihopQA and Musique, potentially associated with retrieval difficulty on Musique-Normal in (Yue et al., 2024).
+These results highlight the importance of both precise retrieval and effective long-context construction for achieving consistent gains across various multi-hop reasoning and long-context datasets.
+Integration with Multi-step Generation Method.
+Integrating MacRAG with the LongRAG (E&F) generation method across multiple LLMs (Llama-3.1-8B, Gemini-1.5-pro, and GPT-4o) yields substantial performance improvements, as shown in Table 1.
+This underscores MacRAG’s versatility and robustness as a retrieval approach when paired with the LongRAG generation method.
+By constructing an effective long context, MacRAG demonstrates advantages in multi-step generation with LongRAG (E&F), producing consistent and meaningful gains across three datasets and three LLMs.
+In contrast, it is not straightforward to extend RAPTOR in combination with LongRAG (E&F).
+As further illustrated in Table 1, the performance gains for Gemini-1.5-pro and GPT-4o indicate that pairing MacRAG with more powerful LLMs can achieve even greater improvements, despite already high F1-scores obtained by other frameworks under these settings.
+We further observe that summarization during indexing improves retrieval quality.
+Our preliminary experiments show a 7.2% gain in precision when using summaries instead of raw slices.
+Robustness across Generation Methods.
+In addition to MacRAG’s robust gains in both vanilla single-step (R&B) and multi-step generation (E&F), MacRAG consistently shows strong advantages across the four metrics listed in Table 2, covering all seven generation modes in Table 4.
+These substantial improvements across various datasets, metrics, and LLMs demonstrate the effectiveness of MacRAG’s long-context retrieval in enhancing both precision and recall for RAG.
+Notably, MacRAG achieves significant gains on the Musique datasets, which Yue et al.
+(2024) identifies as having challenging retrieval conditions.
+Beyond the average gains in these four metrics, Figure 2 further confirms MacRAG’s robust benefits in all seven single-/multi-step generation schemes in Table 4 in Appendix A.1.
+Finally, regardless of the choice of $k_{2}$ or reranker, MacRAG maintains substantial advantages in all test settings in Table 5.
+Enhanced Gains with Stronger Models.
+When applied to the more powerful Gemini-1.5-pro and GPT-4o models, MacRAG showcases its capacity to further improve performance by optimizing the retrieval process and building more relevant, informative contexts.
+Its use of hierarchical chunking and slicing, coupled with adaptive propagation and merging, maintains high precision while expanding coverage.
+This leads to more efficient handling of long documents, reduced computational overhead, and higher-quality generated answers overall.
+Ablation Study.
+Table 3 presents an ablation study demonstrating the effectiveness of MacRAG’s key components: Propagation and Merging, Scaling-up, and Hierarchical Slicing Retrieval.
+Each component contributes to cumulative performance gains.
+The combination of propagation and merging with the scaling-up mechanism increases the coverage of relevant contexts for multi-hop reasoning tasks.
+Figure 3 further illustrates the benefits of scaling up promising candidates with $\alpha=[2,3,4]$, showing how this approach effectively balances context lengths while incorporating additional relevant candidates.
+We observe that performance gains remain consistent across different $k_{2}$ values and rerankers, indicating robustness.
+Although the parameters $(k_{1},k_{2},\alpha,h)$ are fixed, MacRAG adaptively merges nearby slices depending on partial relevance signals.
+Removing this adaptivity leads to a 5% F1 drop, as shown in our ablations.
+4.3 Generation Schemes and Input Lengths
+Figure 4: Cumulative LLM’s input context lengths for single/multi-step generation methods in Table 4.
+While multi-step generation methods (e.g., LongRAG’s Full_E&F) improve answer quality, they often increase cumulative input length.
+Figure 4 compares cumulative input lengths for single and multi-step generation methods in Table 4, given single retrieval with specific settings, highlighting that R&L, Full_Ext, and Full_E&F accumulate the most overhead by relying on entire documents.
+In contrast, MacRAG’s final $k_{2}$ chunks, though potentially larger than basic R&B chunks, preserve bounded context well below full-document with long lengths.
+As Figure 2 and Table 3 illustrate, MacRAG consistently outperforms baselines under both R&B_Ext and R&B_E&F settings.
+Due to MacRAG’s advantage in effective and stable long context construction, even our intermediate R&B_E&F variant demonstrates significant improvements as shown in Table 3.
+For example, on 2WikiMultihopQA (GPT-4o), MacRAG (R&B_E&F) achieves an F1 of 72.43 versus LongRAG (Full_E&F) at 65.89 (+9.93% relative gain), while using 8.19% less context.
+On Musique, the intermediate variant of MacRAG (R&B_E&F) reaches 45.84 compared to LongRAG (Full_E&F)’s 43.83 (+4.58% relative) with a 45.87% reduction in context length.
+This highlights MacRAG’s efficiency in evidence gathering and context construction.
+4.4 Efficiency and Latency Analysis
+MacRAG maintains sub-second efficiency despite constructing richer multi-scale contexts.
+On average, retrieval and reranking take 0.23 seconds per query, which is 38% faster than RAPTOR’s 0.37 seconds.
+With Llama 3.1 8B, total inference latency is 0.99 seconds (0.23s retrieval and 0.76s generation), compared to 0.80 seconds for RAPTOR (0.37s and 0.43s, respectively).
+This slight increase in generation time is offset by improved retrieval precision and higher answer quality on R&B queries.
+This efficiency is enabled by MacRAG’s lightweight, index-based context construction and shared reranking modules, which preserve the same level of computational complexity as baseline retrieval pipelines.
+MacRAG introduces a one-time offline summarization step during indexing, similar to RAPTOR and SIRERAG.
+This step runs in parallel with standard RAG indexing, adding no runtime latency and supporting efficient index maintenance.
+As shown in Table 3, Figure 4, and Subsection 4.3, MacRAG achieves a better trade-off than LongRAG between input context length, answer quality, and overall computational cost.
+5 Conclusion
+In this work, we introduced Multi-scale Adaptive Context RAG (MacRAG), a novel framework designed to address key trade-offs in RAG involving precision, coverage, and computational efficiency for long-context question answering and multi-hop reasoning.
+By dynamically constructing query-adaptive long contexts through real-time, multi-scale retrieval, MacRAG demonstrably improves both precision and recall while maintaining operational efficiency.
+Our extensive experiments validated MacRAG’s advantages across diverse datasets and both single- and multi-step generation paradigms, showcasing consistent and notable performance gains.
+These findings establish MacRAG as a potent and efficient core module for advancing next-generation iterative and agentic RAG systems, with significant potential for robust enterprise-level applications.
+Limitations
+While MacRAG demonstrates strong performance with its modular and efficient design, several avenues offer potential for further extension.
+The current offline summarization and indexing strategy, though effective and its costs amortized, could benefit from adaptive or online variants to better handle rapidly evolving corpora.
+Furthermore, while the use of fixed hyperparameters such as chunk size and expansion scale has shown robustness, dynamically adjusting these parameters based on query complexity presents a promising research direction.
+Finally, tighter integration of MacRAG with sophisticated agentic or multi-round retrieval workflows could further enhance its applicability in highly interactive and real-time RAG scenarios.
+Appendix A Appendix
+A.1 Generation Modes
+Table 4 summarizes different generation modes.
+In addition to MacRAG’s robust gains in both vanilla single-step (R&B) and multi-step generation (E&F), MacRAG consistently shows strong advantages across the four metrics listed in Table 2, covering all seven generation modes in Table 4.
+These substantial improvements across various datasets, metrics, and LLMs demonstrate the effectiveness of MacRAG’s long-context retrieval in enhancing both precision and recall for RAG.
+Notably, MacRAG achieves significant gains on the Musique datasets, which Yue et al.
+(2024) identifies as having challenging retrieval conditions.
+Beyond the average gains in these four metrics, Figure 2 further confirms MacRAG’s robust benefits in all seven single-/multi-step generation schemes in Table 4 in Appendix A.1.
+Finally, regardless of the choice of $k_{2}$ or reranker, MacRAG maintains substantial advantages in all test settings in Table 5.
+Mode Description
+R&B Single-step generation using retrieved top-$k_{2}$ chunks, which is the vanilla basic version.
+R&L Single-step generation using complete documents from top-$k_{2}$ chunks. Prone to the “Lost in the Middle” effect and inefficient for long documents.
+Full_Ext Multi-step generation: (1) LLM extracts query-relevant content from full documents of top-$k_{2}$ chunks, (2) generates answer using both chunks and extracted content.
+Fil Filtering re-ranked top-$k_{2}$ chunks, e.g., via LLM. Improves precision but may lose borderline info.
+Full_E&F Multi-step generation combining Full_Ext’s extracted content and Fil’s filtered chunks. Balances coverage and precision, but is expensive for long documents.
+R&B_Ext Multi-step generation with content extraction limited to top-$k_{2}$ chunks, reducing computational overhead.
+R&B_E&F Multi-step generation combining R&B_Ext’s extracted content and Fil’s filtered chunks. Optimizes precision-recall trade-off w/o processing the full document.
+Table 4: Seven different generation methods with single-step and multi-step generation schemes using retrieved top-$k_{2}$ chunks.
+A.2 Benchmark Datasets and Challenges
+The LongBench variants of HotpotQA, 2WikimultihopQA, and Musique datasets have become the standard evaluation suite for hierarchical retrieval research, including recent systems such as SIRERAG Zhang et al.
 (2025), RAPTOR Sarthi et al.
@@
 (2024).
-Their LongBench variants specifically intensify the “Lost in the Middle" phenomenon by introducing near-duplicate and tangential passages, creating challenging retrieval environments where critical connecting information becomes obscured by surrounding context.
-Each dataset introduces distinct multi-hop reasoning challenges that systematically test different aspects of retrieval capability: HotpotQA features bridging and comparison queries requiring evidence synthesis across multiple paragraphs; 2WikiMultihopQA expands partial excerpts to complete articles, significantly increasing the risk that essential connecting information becomes buried within expansive context; and Musique embeds nested sub-questions within tangential sections, challenging systems to identify and connect scattered bridging facts across document boundaries.
+They specifically intensify the “Lost in the Middle" phenomenon by introducing near-duplicate and tangential passages, creating challenging retrieval environments where critical connecting information becomes obscured by surrounding context.
+Each dataset introduces distinct multi-hop reasoning challenges that systematically test different aspects of retrieval capability: HotpotQA features bridging and comparison queries requiring evidence synthesis across multiple paragraphs.
+2WikiMultihopQA expands partial excerpts to complete articles, significantly increasing the risk that essential connecting information becomes buried within expansive context.
+Musique embeds nested sub-questions within tangential sections, challenging systems to identify and connect scattered bridging facts across document boundaries.
 These characteristics directly test a retrieval system’s ability to handle fragmented evidence, reconcile overlapping information, and identify critical bridging facts despite contextual noise (Yue et al., 2024; Leng et al., 2024), precisely the challenges that MacRAG’s multi-scale architecture addresses through hierarchical indexing and adaptive retrieval.
-Models and Evaluation Metrics:
+A.3 Target Baseline Methods
 Based on SIRERAG’s comparative analysis (Table 4 in Zhang et al.
@@
 (2025).
-Among these top-performing systems, we selected RAPTOR and LongRAG due to their better balance between performance and computational efficiency, offering lower computational costs and easier maintenance compared to SIRERAG’s higher overhead.
-We conducted experiments using three prominent long-context LLMs: two widely adopted proprietary models (GPT-4o and Gemini-1.5-pro) and one popular open-source model (Llama-3.1-8B-instruct).
-This diverse selection enables assessment of MacRAG’s generalizability across different model architectures and parameter scales.
-We measure performance through four complementary metrics: Exact Match, F1-score, Precision, and Recall, providing a multidimensional view of answer quality.
-We additionally track average context length to quantify MacRAG’s efficiency in reducing input redundancy while maintaining or improving information coverage.
-To ensure fair comparison across all methods (MacRAG, LongRAG, and RAPTOR), we basically applied the same re-ranking model (marco-miniLM) for all RAG methods.
-Comparative Framework and Hyperparameters:
-To ensure rigorous comparative analysis across all methods (MacRAG, LongRAG, and RAPTOR), we used the same re-ranking model (marco-miniLM) for all RAG methods.
-We also adopted the optimal hyperparameter configuration from LongRAG Zhao et al.
-(2024), and used identical hyperparameters across all methods: $k_{1}=100$ chunks retrieved initially, $k_{2}=7$ chunks selected after reranking, except in ablation studies.
-We maintain identical prompt templates across all methods to eliminate prompt engineering as a confounding variable.
-For MacRAG-specific parameters, we varied the scale-up factor $\alpha\in\{1,4\}$ and the neighbor hop count $h\in\{0,1\}$, except during ablation studies.
-The ablation study and performance analysis in Section 4.2 demonstrate both robustness and the soundness of the architecture design of MacRAG, enhancing multi-hop reasoning capability.
-This carefully controlled experimental framework ensures that observed performance differences can be attributed specifically to MacRAG’s architectural innovations, particularly hierarchical document representation and adaptive, multi-scale retrieval rather than variations in hyperparameter optimization or underlying retrieval technologies.
-By maintaining consistency while isolating architectural factors, we can precisely measure the impact of multi-scale retrieval on multi-hop reasoning performance.
-4.2 Main Results
+Among these top-performing systems, we selected RAPTOR and LongRAG because they balance performance and computational efficiency better, offering lower computational costs and easier maintenance than SIRERAG’s higher overhead.
+A.4 Adaptive Retrieval Mechanism Analysis
+A key strength of MacRAG lies in its adaptive retrieval mechanism, which dynamically constructs query-specific contexts despite using fixed common hyperparameters ($k_{1}$, $k_{2}$) and MacRAG’s $\alpha$, and $h$.
+This adaptivity operates through several complementary mechanisms:
+Query-dependent Adaptive Context Expansion:
+For complex multi-hop queries, MacRAG naturally extends its context boundaries to encompass connecting information.
+For simpler queries, expansion remains more focused, selectively around the most promising content regions, but avoiding unnecessary content.
+Our ablation studies quantify the impact of this adaptivity.
+When the adaptivity of query-adaptive expansion of neighborhoods is removed by eliminating the propagation and merging steps (Table 3), performance drops by up to around 5% F1-score on datasets.
+This degradation is most pronounced on Musique, where connecting multiple scattered pieces of evidence is crucial.
+Table 3 additionally reveals another important aspect of query-adaptive expansion of borderline candidates "Scaling-up", and if we eliminate its corresponding "Scaling-up" component, then performance drops up to around 6% F1-score on datasets.
+This aligns with our expectation that adaptive expansion becomes increasingly valuable as query complexity grows.
+A.5 Preprocessing Efficiency of MacRAG
+Chunk Size and Token Budget:
+MacRAG uses original chunk sizes of approximately 400 tokens (around 1500 characters), which fall within the 200–500 token range used in prior work such as LongRAG even for re-ranking on retrieved chunks.
+The average sizes of summarized chunks are around 800-1000 characters.
+We applied slices into summarized chunks for initial retrieval to improve precision further with two slice sizes around 450 characters with 300 character overlaps or 600 characters with 450 character overlaps.
+These sizes allow MacRAG to maintain sub-second retrieval and reranking times while keeping the context structure manageable for downstream language models.
+Offline Summarization Cost:
+MacRAG includes a one-time summarization step during indexing, which does not affect query-time latency.
+This step is comparable in cost to offline embedding computations in standard RAG pipelines.
+All runtime efficiency measurements reported here exclude this indexing cost, as it is amortized over the use of the document collection.
+A.6 Latency Comparison between MacRAG and RAPTOR
+Retrieval and Reranking Latency:
+MacRAG maintains sub-second latency for retrieval and reranking across all evaluated datasets.
+Using dense retrieval with $k_{1}=100$ followed by cross-encoder reranking with $k_{2}=7$, the measured latency is 0.23 seconds on HotpotQA, 0.22 seconds on 2WikimultihopQA, and 0.24 seconds on Musique.
+In contrast, RAPTOR reports 0.43s, 0.24s, and 0.44s on the same datasets, averaging 0.37s.
+This represents a 38% speedup on average, even though both methods operate on the same top-100 retrieved chunks and use the same reranker (marco-miniLM).
+MacRAG achieves this by leveraging index-based merging after reranking, which reuses precomputed relevance scores and adds only a few milliseconds per query.
+Generation Time and Trade-offs:
+While retrieval and reranking are efficient, generation time varies depending on the context length produced by each method.
+MacRAG’s adaptive multi-scale expansion typically results in longer but more relevant contexts.
+With Llama 3.1 8B, the average generation time with R&B is 0.76 seconds, compared to 0.43 seconds for RAPTOR with R&B.
+This increase reflects MacRAG’s strategy of incorporating broader evidence to improve answer accuracy.
+Users seeking faster inference can reduce the expansion parameter $\alpha$ to trade off coverage for speed, though at the cost of some performance.
+Total Latency Comparison:
+Combining retrieval, reranking, and generation with R&B, the total end-to-end latency of MacRAG is 0.99 seconds (0.23s + 0.76s), compared to RAPTOR’s 0.80 seconds (0.37s + 0.43s).
+While MacRAG incurs a small additional cost, it consistently yields higher accuracy across multiple datasets and models, as shown in Table 1.
+A.7 Efficiency Comparison between MacRAG and LongRAG
+MacRAG and LongRAG both rely on index-based hierarchical retrieval and reranking over the same number of $k_{1}$ chunks.
+As a result, their retrieval and reranking steps show similar efficiency and latency.
+However, for the generation step, according to Table 3, Figure 4, and Subsection 4.3, MacRAG achieves improved efficiency compared to LongRAG in terms of the trade-off between final answer quality, input context length, and overall computational cost.
+A.8 Discussion on Effectiveness of MacRAG’s Retrieval and Constructed Long Context
+Enhanced Precision and Coverage:  MacRAG employs a hierarchical multi-scale strategy that begins with fine-grained slice retrieval and chunk-level re-ranking for precise identification of relevant content, then systematically incorporates broader document-level context through $h$-hop neighbor expansions.
+To preserve coherence without overwhelming the LLM, it strategically up-scales via the $(\alpha\times k_{2})$ factor, capturing borderline yet crucial segments.
+Rather than relying on compressed text or entire documents, MacRAG focuses on real-time construction of effective long contexts from promising original chunks while enforcing an upper bound on context length.
+By removing irrelevant portions from documents and maintaining a continuous, coherent subset of text, MacRAG avoids excessive token consumption, mitigates the “lost in the middle" phenomenon, and minimizes hallucinations, ultimately ensuring high recall for complex multi-hop queries.
+A.9 Discussion on Single, Multi-Step, Iterative, and Agentic Generation with MacRAG
+MacRAG’s modular design seamlessly supports single and multi-step generation, iterative retrieval, and agentic pipelines.
+By combining its core multi-scale retrieval with standard generation methods, MacRAG dynamically adapts context formation and refinement across multiple rounds without overwhelming the LLM.
+In single and multi-step modes (e.g., LongRAG Zhao et al.
+(2024) and standard multi-step QA in Section 3.3), MacRAG selects and assembles relevant text from large corpora while minimizing noise and retaining crucial connections, thereby improving both recall and precision for multi-hop reasoning.
+For iterative scenarios, such as IterDRAG Yue et al.
+(2024) or chain-of-RAG Wang et al.
+(2025), MacRAG can support efficient updates on the evidence set via $\mathcal{C}_{t}\leftarrow\mathrm{Merge}\Bigl{(}\mathcal{C}_{t-1},\,\mathrm{Retrieval}(q_{t})\Bigr{)}$ to unify old and newly retrieved content, promoting consistent coverage of bridging facts while discarding irrelevant material.
+In agentic RAG settings Asai et al.
+(2024); Jeong et al.
+(2024); Chen et al.
+(2025), MacRAG likewise prevents context explosion by focusing specifically on segments required for each action, thereby improving precision and recall over multiple steps.
+Our experiments (Section 4) confirm that integrating MacRAG with single and multi-step generation methods, potential core modules for multi-round advanced systems, consistently enhances complex multi-hop reasoning tasks and reduces error rates in long-context RAG.
+A.10 Discussion on Graph-Enhanced MacRAG
+A promising direction is extending MacRAG with a two-stage reranking strategy incorporating graph-based knowledge structures.
+After initial chunk-level reranking, MacRAG could perform a second reranking phase by expanding top candidates through both local index-based extensions and pre-constructed graph neighbors, enabling efficient coverage of both local and global relationships.
+This approach would leverage MacRAG’s bounded context guarantees at each phase while allowing controlled exploration of knowledge-guided connections, managing computational efficiency through bounded candidate sets during reranking.
+By applying MacRAG’s document-oriented indexing or relationship-based thresholding to merge expanded candidates, this strategy would enhance retrieval quality for complex queries requiring both precise local context and broad knowledge integration.
 Model $\mathbf{k_{2}}$ Reranker R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_E&F Average
@@
 (2024).
-($k_{2}$=7, marco-miniLM) R&B R&L Full_Ext Fil Full_E&F R&B_Ext R&B_E&F Average
-HotpotQA
-LongRAG 65.46 69.40 66.60 63.32 66.20 66.08 66.09 66.16
-MacRAG 67.15 (+1.69$\uparrow$) 69.55 (+0.15$\uparrow$) 69.00 (+2.40$\uparrow$) 65.44 (+2.12$\uparrow$) 68.52 (+2.32$\uparrow$) 66.36 (+0.28$\uparrow$) 66.53 (+0.44$\uparrow$) 67.51 (+1.35$\uparrow$, +2.04$\uparrow$)
-- Propagation&Merging 64.17 (-1.29$\downarrow$) 68.91 (-0.49$\downarrow$) 67.82 (+1.22$\uparrow$) 68.06 (+4.74$\uparrow$) 68.44 (+2.24$\uparrow$) 63.77 (-2.31$\downarrow$) 65.57 (-0.52$\downarrow$) 66.68 (+0.52$\uparrow$, +0.79$\uparrow$)
-- Scaling up 64.12 (-1.34$\downarrow$) 67.57 (-1.83$\downarrow$) 67.33 (+0.73$\uparrow$) 64.41 (+1.09$\uparrow$) 67.26 (+1.06$\uparrow$) 64.22 (-1.86$\downarrow$) 63.84 (-2.44$\downarrow$) 65.54 (-0.62$\downarrow$, -0.94$\downarrow$)
-2WikimultihopQA
-LongRAG 59.97 62.37 65.35 60.68 65.89 62.08 62.47 62.69
-MacRAG 59.00 (-0.97$\downarrow$) 64.87 (+2.50$\uparrow$) 68.97 (+3.62$\uparrow$) 68.20 (+7.52$\uparrow$) 73.19 (+7.30$\uparrow$) 66.50 (+4.42$\uparrow$) 71.40 (+8.93$\uparrow$) 67.45 (+4.76$\uparrow$, +7.59$\uparrow$)
-- Propagation&Merging 59.67 (-0.30$\downarrow$) 65.06 (+2.69$\uparrow$) 68.48 (+3.13$\uparrow$) 67.57 (+6.89$\uparrow$) 72.20 (+6.31$\uparrow$) 65.13 (+3.05$\uparrow$) 72.43 (+9.96$\uparrow$) 67.22 (+4.53$\uparrow$, +7.23$\uparrow$)
-- Scaling up 59.00 (-0.97$\downarrow$) 61.89 (-0.48$\downarrow$) 67.63 (+2.28$\uparrow$) 60.41 (-0.27$\downarrow$) 67.03 (+1.14$\uparrow$) 66.37 (+4.29$\uparrow$) 65.77 (+3.30$\uparrow$) 64.01 (+1.32$\uparrow$, +2.11$\uparrow$)
-Musique
-LongRAG 38.98 41.90 43.64 37.72 43.83 42.00 41.97 41.43
-MacRAG 44.76 (+5.78$\uparrow$) 45.74 (+3.84$\uparrow$) 47.42 (+3.78$\uparrow$) 47.80 (+10.08$\uparrow$) 50.09 (+6.26$\uparrow$) 48.57 (+6.57$\uparrow$) 51.00 (+9.03$\uparrow$) 47.91 (+6.48$\uparrow$, +15.63$\uparrow$)
-- Propagation&Merging 41.61 (+2.63$\uparrow$) 45.43 (+3.53$\uparrow$) 48.00 (+4.36$\uparrow$) 47.48 (+9.76$\uparrow$) 52.26 (+8.43$\uparrow$) 40.09 (-1.91$\downarrow$) 45.84 (+3.87$\uparrow$) 45.82 (+4.39$\uparrow$, +10.58$\uparrow$)
-- Scaling up 41.65 (+2.67$\uparrow$) 45.87 (+3.97$\uparrow$) 46.80 (+3.16$\uparrow$) 41.83 (+4.11$\uparrow$) 46.59 (+2.76$\uparrow$) 39.40 (-2.60$\downarrow$) 39.63 (-2.34$\downarrow$) 43.11 (+1.68$\uparrow$, +4.04$\uparrow$)
-Table 5: Ablation Study of MacRAG with GPT-4o and F1-score. The experiments conducted with the same $k_{1}$, $k_{2}$ and reranker.
-Figure 3: Performance trends across datasets for scale factors ($\alpha$).
-Comparison with RAPTOR:  While RAPTOR Sarthi et al.
-(2024), a summarization and hierarchical clustering-based approach, shows gains on standard datasets, it struggles with retrieval quality on LongBench versions Bai et al.
-(2023) as shown in Table 2.
-Specifically, RAPTOR with Llama-3.1-8B shows performance decreases in 2WikimultihopQA and Musique, potentially associated with retrieval difficulty on Musique-Normal in (Yue et al., 2024).
-These results highlight the importance of both precise retrieval and effective long context construction for achieving consistent gains across various multi-hop reasoning and long-context datasets.
-Integration with Multi-step Generation Method:  Integrating MacRAG with the LongRAG (E&F) generation method across multiple LLMs (Llama-3.1-8B, Gemini-1.5-pro, and GPT-4o) yields substantial performance improvements, as shown in Table 2.
-This underscores MacRAG’s versatility and robustness as a retrieval approach when paired with the LongRAG generation method.
-By constructing an effective long context, MacRAG demonstrates advantages in multi-step generation with LongRAG (E&F), producing consistent and meaningful gains across three datasets and three LLMs.
-In contrast, it is not straightforward to extend RAPTOR in combination with LongRAG (E&F).
-As further illustrated in Table 2, the performance gains for Gemini-1.5-pro and GPT-4o indicate that pairing MacRAG with more powerful LLMs can achieve even greater improvements, despite already high F1-scores obtained by other frameworks under these settings.
-We further observe that summarization during indexing improves retrieval quality.
-Our preliminary experiments show a 7.2% gain in precision when using summaries instead of raw slices.
-Consistent Gains across Generation Methods:  In addition to MacRAG’s robust gains in both vanilla single-step (R&B) and multi-step generation (E&F), MacRAG consistently shows strong advantages across the four metrics listed in Table 3, covering all seven generation modes in Table 1.
-These substantial improvements across various datasets, metrics, and LLMs demonstrate the effectiveness of MacRAG’s long-context retrieval in enhancing both precision and recall for RAG.
-Notably, MacRAG achieves significant gains on the Musique datasets, which Yue et al.
-(2024) identifies as having challenging retrieval conditions.
-Beyond the average gains in these four metrics, Figure 2 further confirms MacRAG’s robust benefits in all seven single-/multi-step generation schemes in Table 1.
-Finally, regardless of the choice of $k_{2}$ or reranker, MacRAG maintains substantial advantages in all test settings in Table 4.
-Enhanced Gains with Stronger Models: When applied to the more powerful Gemini-1.5-pro and GPT-4o models, MacRAG showcases its capacity to further improve performance by optimizing the retrieval process and building more relevant, informative contexts.
-Its use of hierarchical chunking and slicing, coupled with adaptive propagation and merging, maintains high precision while expanding coverage.
-This leads to more efficient handling of long documents, reduced computational overhead, and higher-quality generated answers overall.
-Ablation Study:  Table 5 presents an ablation study demonstrating the effectiveness of MacRAG’s key components: Propagation and Merging, Scaling-up, and Hierarchical Slicing Retrieval.
-Each component contributes to cumulative performance gains.
-The combination of propagation and merging with the scaling-up mechanism increases the coverage of relevant contexts for multi-hop reasoning tasks.
-Figure 3 further illustrates the benefits of scaling up promising candidates with $\alpha=[2,3,4]$, showing how this approach effectively balances context lengths while incorporating additional relevant candidates.
-We observe that performance gains remain consistent across different $k_{2}$ values and rerankers, indicating robustness.
-Although the parameters $(k_{1},k_{2},\alpha,h)$ are fixed, MacRAG adaptively merges nearby slices depending on partial relevance signals.
-Removing this adaptivity leads to a 5% F1 drop, as shown in our ablations.
-4.3 Generation Schemes and Input Lengths
-Figure 4: Cumulative LLM’s input context lengths across steps for single/multi-step generation methods in Table 1.
-While multi-step generation can improve answer quality, it also increases total input length across retrieval and generation steps for a single round.
-Figure 4 compares cumulative input lengths for single and multi-step generation methods in Table 1, given single retrieval with specific settings, highlighting that R&L, Full_Ext, and Full_E&F accumulate the most overhead by relying on entire documents.
-In contrast, MacRAG’s final $k_{2}$ chunks, though potentially larger than basic R&B chunks, preserve bounded context well below full-document with long lengths.
-As Figure 2 and Table 5 illustrate, MacRAG consistently outperforms baselines under both R&B_Ext and R&B_E&F settings.
-Due to MacRAG’s advantage in effective and stable long context construction, even our intermediate R&B_E&F variant demonstrates significant improvements as shown in Table 5.
-While we do not report retrieval recall directly, MacRAG achieves competitive performance with reduced input length, suggesting effective evidence gathering through its hierarchical expansion.For instance, on 2WikimultihopQA, MacRAG (R&B_E&F) achieves an F1-score of 72.43 versus LongRAG (Full_E&F) at 65.89 (+6.54 absolute, +9.93% relative), using 8.19% less context.
-On Musique, the intermediate variant of MacRAG (R&B_E&F) reaches 45.84 compared to LongRAG (Full_E&F)’s 43.83 (+2.01 absolute, +4.58% relative) with a 45.87% reduction in context length.
-4.4 Efficiency and Latency Analysis
-MacRAG introduces a one-time offline summarization step during indexing like other methods (RAPTOR and SIRERAG).
-The summarization step for MacRAG provides both parallel standard RAG indexing, which does not affect real-time latency, and efficient maintenance.
-Despite constructing richer multi-scale contexts, MacRAG maintains efficient retrieval and reranking, averaging 0.23 seconds across datasets, which is 38% faster than RAPTOR’s 0.37 seconds.
-The total inference time with Llama 3.1 8B is 0.99 seconds for MacRAG (0.23s retrieval + 0.76s generation), compared to 0.80 seconds for RAPTOR (0.37s + 0.43s) on typical multi-hop queries, balancing the added overhead with improved downstream accuracy on R&B.
-During inference, MacRAG’s multi-scale adaptive expansions preserve the same level of computational complexity as baseline sub-second retrieval and reranking, thanks to shared reranking components and lightweight, index-based context construction.
-Moreover, as shown in Table 5, Figure 4, and Subsection 4.3, MacRAG demonstrates improved efficiency over LongRAG in terms of the trade-off between final answer quality, input context length, and overall computational cost.
-5 Conclusion
-In this work, we introduce the Multi-scale Adaptive Context RAG (MacRAG) framework, which addresses the key trade-offs in RAG involving precision, coverage, and computational overhead for long-context question answering and multi-hop reasoning.
-By constructing query-adaptive, effective long contexts in real-time retrieval, MacRAG simultaneously improves precision and recall while maintaining efficiency.
-Experiments demonstrated its advantages with both single- and multi-step generation on multi-hop and large-document datasets, achieving notable performance gains.
-These findings underscore MacRAG’s potential as a core module for iterative and agentic RAG systems, offering substantial benefits for enterprise-level applications.
-Acknowledgments
-We are grateful to Prof. Xifeng Yan for his valuable feedback and suggestions throughout this research.
-Appendix A Appendix
-A.1 Adaptive Retrieval Mechanism Analysis
-A key strength of MacRAG lies in its adaptive retrieval mechanism, which dynamically constructs query-specific contexts despite using fixed common hyperparameters ($k_{1}$, $k_{2}$) and MacRAG’s $\alpha$, and $h$.
-This adaptivity operates through several complementary mechanisms:
-Query-dependent Adaptive Context Expansion:
-For complex multi-hop queries, MacRAG naturally extends its context boundaries to encompass connecting information.
-For simpler queries, expansion remains more focused, selectively around the most promising content regions, but avoiding unnecessary content.
-Our ablation studies quantify the impact of this adaptivity.
-When the adaptivity of query-adaptive expansion of neighborhoods is removed by eliminating the propagation and merging steps (Table 5), performance drops by up to around 5% F1-score on datasets.
-This degradation is most pronounced on Musique, where connecting multiple scattered pieces of evidence is crucial.
-Table 5 additionally reveals another important aspect of query-adaptive expansion of borderline candidates "Scaling-up", and if we eliminate its corresponding "Scaling-up" component, then performance drops up to around 6% F1-score on datasets.
-This aligns with our expectation that adaptive expansion becomes increasingly valuable as query complexity grows.
-A.2 Preprocessing Efficiency of MacRAG
-Chunk Size and Token Budget:
-All methods in our comparison use chunk sizes of approximately 400 tokens (around 1500 characters), which fall within the 200–500 token range used in prior work such as LongRAG.
-These sizes allow MacRAG to maintain sub-second retrieval and reranking times while keeping the context structure manageable for downstream language models.
-Offline Summarization Cost:
-MacRAG includes a one-time summarization step during indexing, which does not affect query-time latency.
-This step is comparable in cost to offline embedding computations in standard RAG pipelines.
-All runtime efficiency measurements reported here exclude this indexing cost, as it is amortized over the use of the document collection.
-A.3 Latency Comparison between MacRAG and RAPTOR
-Retrieval and Reranking Latency:
-MacRAG maintains sub-second latency for retrieval and reranking across all evaluated datasets.
-Using dense retrieval with $k_{1}=100$ followed by cross-encoder reranking with $k_{2}=7$, the measured latency is 0.23 seconds on HotpotQA, 0.22 seconds on 2WikimultihopQA, and 0.24 seconds on Musique.
-In contrast, RAPTOR reports 0.43s, 0.24s, and 0.44s on the same datasets, averaging 0.37s.
-This represents a 38% speedup on average, even though both methods operate on the same top-100 retrieved chunks and use the same reranker (marco-miniLM).
-MacRAG achieves this by leveraging index-based merging after reranking, which reuses precomputed relevance scores and adds only a few milliseconds per query.
-Generation Time and Trade-offs:
-While retrieval and reranking are efficient, generation time varies depending on the context length produced by each method.
-MacRAG’s adaptive multi-scale expansion typically results in longer but more relevant contexts.
-With Llama 3.1 8B, the average generation time with R&B is 0.76 seconds, compared to 0.43 seconds for RAPTOR with R&B.
-This increase reflects MacRAG’s strategy of incorporating broader evidence to improve answer accuracy.
-Users seeking faster inference can reduce the expansion parameter $\alpha$ to trade off coverage for speed, though at the cost of some performance.
-Total Latency Comparison:
-Combining retrieval, reranking, and generation with R&B, the total end-to-end latency of MacRAG is 0.99 seconds (0.23s + 0.76s), compared to RAPTOR’s 0.80 seconds (0.37s + 0.43s).
-While MacRAG incurs a small additional cost, it consistently yields higher accuracy across multiple datasets and models, as shown in Table 2.
-A.4 Efficiency Comparison between MacRAG and LongRAG
-MacRAG and LongRAG both rely on index-based hierarchical retrieval and reranking over the same number of $k_{1}$ chunks.
-As a result, their retrieval and reranking steps show similar efficiency and latency.
-However, for the generation step, according to Table 5, Figure 4, and Subsection 4.3, MacRAG achieves improved efficiency compared to LongRAG in terms of the trade-off between final answer quality, input context length, and overall computational cost.
-A.5 Discussion on Effectiveness of MacRAG’s Retrieval and Constructed Long Context
-Enhanced Precision and Coverage:  MacRAG employs a hierarchical multi-scale strategy that begins with fine-grained slice retrieval and chunk-level re-ranking for precise identification of relevant content, then systematically incorporates broader document-level context through $h$-hop neighbor expansions.
-To preserve coherence without overwhelming the LLM, it strategically up-scales via the $(\alpha\times k_{2})$ factor, capturing borderline yet crucial segments.
-Rather than relying on compressed text or entire documents, MacRAG focuses on real-time construction of effective long contexts from promising original chunks while enforcing an upper bound on context length.
-By removing irrelevant portions from documents and maintaining a continuous, coherent subset of text, MacRAG avoids excessive token consumption, mitigates the “lost in the middle" phenomenon, and minimizes hallucinations, ultimately ensuring high recall for complex multi-hop queries.
-A.6 Discussion on Single, Multi-Step, Iterative, and Agentic Generation with MacRAG
-MacRAG’s modular design seamlessly supports single and multi-step generation, iterative retrieval, and agentic pipelines.
-By combining its core multi-scale retrieval with standard generation methods, MacRAG dynamically adapts context formation and refinement across multiple rounds without overwhelming the LLM.
-In single and multi-step modes (e.g., LongRAG Zhao et al.
-(2024) and standard multi-step QA in Section 3.3), MacRAG selects and assembles relevant text from large corpora while minimizing noise and retaining crucial connections, thereby improving both recall and precision for multi-hop reasoning.
-For iterative scenarios, such as IterDRAG Yue et al.
-(2024) or chain-of-RAG Wang et al.
-(2025), MacRAG can support efficient updates on the evidence set via $\mathcal{C}_{t}\leftarrow\mathrm{Merge}\Bigl{(}\mathcal{C}_{t-1},\,\mathrm{Retrieval}(q_{t})\Bigr{)}$ to unify old and newly retrieved content, promoting consistent coverage of bridging facts while discarding irrelevant material.
-In agentic RAG settings Asai et al.
-(2024); Jeong et al.
-(2024); Chen et al.
-(2025), MacRAG likewise prevents context explosion by focusing specifically on segments required for each action, thereby improving precision and recall over multiple steps.
-Our experiments (Section 4) confirm that integrating MacRAG with single and multi-step generation methods, potential core modules for multi-round advanced systems, consistently enhances complex multi-hop reasoning tasks and reduces error rates in long-context RAG.
-A.7 Discussion on Graph-Enhanced MacRAG
-A promising direction is extending MacRAG with a two-stage reranking strategy incorporating graph-based knowledge structures.
-After initial chunk-level reranking, MacRAG could perform a second reranking phase by expanding top candidates through both local index-based extensions and pre-constructed graph neighbors, enabling efficient coverage of both local and global relationships.
-This approach would leverage MacRAG’s bounded context guarantees at each phase while allowing controlled exploration of knowledge-guided connections, managing computational efficiency through bounded candidate sets during reranking.
-By applying MacRAG’s document-oriented indexing or relationship-based thresholding to merge expanded candidates, this strategy would enhance retrieval quality for complex queries requiring both precise local context and broad knowledge integration.
 Figure 5: Performances of LongRAG and MacRAG regarding the fours metrics (Exact Match, F1-score, Precision, Recall) for three datasets (HotpotQA, 2WikimultihopQA, and Musique) and two LLMs (Gemini-1.5-pro and GPT-4o).